### v1

# Matrix systems, algebras, and open maps

Abstract.
Every state on the algebra $\mathrm{M}_{n}$ of complex $n\times n$ matrices restricts to a state on any matrix system. Whereas the restriction to a matrix system is generally not open, we prove that the restriction to every *-subalgebra of $\mathrm{M}_{n}$ is open. This simplifies topology problems in matrix theory and quantum information theory.
Key words and phrases:
open map, numerical range, matrix system, quantum state
2020 Mathematics Subject Classification:
46A30, 15A60, 47A12, 81P16
In honor of Ilya Matveevich Spitkovsky, for his 70th birthday.
1. Introduction
[…]
$\mathcal{S}(\mathcal{R})$ :=$:=\{\ell\in\mathcal{R}^{\ast}\mid\forall A\in\mathcal{C}(\mathcal{R}):\ell(A)\geq 0,\ell(\mathds{1}_{n})=1\}\,.$
The restriction $\mathcal{S}(\mathrm{M}_{n})\to\mathcal{S}(\mathcal{R})$, $\ell\mapsto\ell|_{\mathcal{R}}$ of states to $\mathcal{R}$ is continuous and affine. Its analytic properties would be perfectly clear if it were not for the openness that fails in Example 1.1. Let $K,L$ be subsets of some Euclidean spaces endowed with their relative topologies [27]. A map $f:K\to L$ is open at $x\in K$ if the image of every neighborhood of $x$ in $K$ is a neighborhood of $f(x)$ in $L$. The map $f$ is open if it is open at every point in $K$.
It is helpful to represent states as matrices. The antilinear isomorphism $\mathcal{R}\to\mathcal{R}^{\ast}$, $A\mapsto\langle A,\,\cdot\,\rangle$ restricts by Lemma 2.2 to the affine isomorphism
$r_{\mathcal{R}}:\mathcal{D}(\mathcal{R})\to\mathcal{S}(\mathcal{R})\,,\quad\rho\mapsto\langle\rho,\,\cdot\,\rangle\,,$
where $\langle A,B\rangle:=\operatorname{tr}(A^{\ast}B)$ is the Frobenius inner product of $A,B\in\mathrm{M}_{n}$,
$\mathcal{H}(\mathcal{R})$ :=$:=\{A\in\mathcal{R}\mid A^{\ast}=A\}\,,$
$\mathcal{C}(\mathcal{R})^{\vee}$ :=$:=\{A\in\mathcal{H}(\mathcal{R})\mid\forall B\in\mathcal{C}(\mathcal{R}):\langle A,B\rangle\geq 0\}\,,$
$\text{and}\quad\mathcal{D}(\mathcal{R})$ :=$:=\{\rho\in\mathcal{C}(\mathcal{R})^{\vee}\mid\operatorname{tr}(\rho)=1\}\,.$
Generalizing a term of von Neumann algebras [10], we refer to the elements of $\mathcal{D}(\mathcal{R})$ as density matrices. The inclusion $\mathcal{C}(\mathcal{R})\subset\mathcal{C}(\mathcal{R})^{\vee}$ can be strict. For example, the density matrix $\operatorname{diag}(-1,5,2)/6$ of the matrix system spanned by $\mathds{1}_{3}$ and $\operatorname{diag}(1,-1,0)$ is indefinite. It is well known, see Rem. 2.8, that the identity $\mathcal{C}(\mathcal{A})=\mathcal{C}(\mathcal{A})^{\vee}$ holds for every *-subalgebra $\mathcal{A}$ of $\mathrm{M}_{n}$.
Example 1.1.
We write block diagonal matrices as direct sums, for instance
$\begin{pmatrix}A&0\\ 0&c\end{pmatrix}=A\oplus c\in\mathrm{M}_{3}\quad\text{for every}\quad A\in\mathrm{M}_{2}\,,\quad c\in\mathbb{C}\cong\mathrm{M}_{1}\,.$
Denoting the imaginary unit by $\mathrm{i}\in\mathbb{C}$ and the Pauli matrices by
$X:=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}\,,\quad Y:=\begin{pmatrix}0&-\mathrm{i}\\ \mathrm{i}&0\end{pmatrix}\,,\quad Z:=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}\,,$
we define the matrix system $\mathcal{R}:=\operatorname{span}_{\mathbb{C}}(\mathds{1}_{3},X\oplus 1,Z\oplus 0)$. The orthogonal projection of $\mathbb{C}^{2}$ onto the line spanned by |$\ket{+}:=(1,1){}^{\mathrm{T}}/\sqrt{2}$ is written $\ket{+}\!\!\bra{+}$ | in Dirac’s notation [7, 11]. The open set $\mathcal{O}:=\{\ell\in\mathcal{S}(\mathrm{M}_{3})\mid\ell(0\oplus 1)>0\}$ contains
$\omega_{\lambda}:=r_{\mathrm{M}_{3}}\big{[}(1-\lambda)\ket{+}\!\!\bra{+}\oplus\lambda\big{]}\,,\quad\lambda\in(0,1]\,.$
So $\mathcal{O}|_{\mathcal{R}}:=\{\ell|_{\mathcal{R}}:\ell\in\mathcal{O}\}$ contains $\omega_{\lambda}|_{\mathcal{R}}$ but none of the restriction $\ell_{\theta}|_{\mathcal{R}}$ of
$\ell_{\theta}:=r_{\mathrm{M}_{3}}\big{[}\frac{1}{2}\big{(}\mathds{1}_{2}+\cos(\theta)X+\sin(\theta)Z\big{)}\oplus 0\big{]}\,,\quad\theta\in(0,2\pi)\,.$
Specifically, $\ell_{\theta}$ has the value $\ell_{\theta}(A_{\theta})=1$ at $A_{\theta}:=\cos(\theta)(X\oplus 1)+\sin(\theta)Z\oplus 0$ and $\ell(A_{\theta})\leq(\cos(\theta)-1)\ell(0\oplus 1)+1$ holds for every $\ell\in\mathcal{S}(\mathrm{M}_{3})$. Hence
$\ell|_{\mathcal{R}}(A_{\theta})-\ell_{\theta}|_{\mathcal{R}}(A_{\theta})\leq(\cos(\theta)-1)\ell(0\oplus 1)<0\,,\quad\ell\in\mathcal{O}\,.$
This shows that $\mathcal{O}|_{\mathcal{R}}$ is not a neighborhood of $\omega_{\lambda}|_{\mathcal{R}}$ as $\lim_{\theta\to 0}\ell_{\theta}|_{\mathcal{R}}=\omega_{\lambda}|_{\mathcal{R}}$. In conclusion, $\mathcal{S}(\mathrm{M}_{3})\to\mathcal{S}(\mathcal{R})$, $\ell\mapsto\ell|_{\mathcal{R}}$ is not open at $\omega_{\lambda}$ for any $\lambda\in(0,1]$.
Asking where $\mathcal{S}(\mathrm{M}_{n})\to\mathcal{S}(\mathcal{R})$, $\ell\mapsto\ell|_{\mathcal{R}}$ is open is the same, by Coro. 2.7 b), as inquiring at which density matrices the orthogonal projection
$\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathcal{R})$ (1.1)
is open. The map (1.1) is defined in (2.2) and (2.3) below. For now suffice it to say that it is a restriction of the orthogonal projection of $\mathrm{M}_{n}$ onto $\mathcal{R}$.
Corey et al. [18] and Leake et al. [29, 30] first studied a problem of numerical ranges closely related to (1.1). The problem (1.1) was studied by Weis [52, 53] and Rodman et al. [46] when $\mathcal{D}(\mathcal{R})$ is replaced with the affinely isomorphic joint numerical range. Numerical ranges are the topic of Sec. 4.
Theorem 1.2.
If $\mathcal{A}$ is a *-subalgebra of $\mathrm{M}_{n}$, then the orthogonal projection $\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathcal{A})$ is open.
Thm. 1.2 can simplify the problem (1.1), and related continuity problems, if $\mathcal{R}$ is included in a *-subalgebra of $\mathrm{M}_{n}$ smaller than $\mathrm{M}_{n}$, as we show in Sec. 8. Examples from quantum information theory are presented in Sec. 9.
Thm. 1.2 is proved in Sec. 6. The main ideas are that $\mathcal{A}$ is a direct sum of full matrix algebras and that $\mathcal{D}(\mathrm{M}_{n})$ is stable and highly symmetric. Thereby, a convex set $K$ is stable if the midpoint map
$K\times K\to K\,,\quad(x,y)\mapsto\frac{1}{2}(x+y)$
is open. A convex set is always (except Rem. 1.3) understood to be included in a Euclidean space. Debs [20] proved the stability of $\mathcal{D}(\mathcal{A})$ for $\mathcal{A}=\mathrm{M}_{n}$, Papadopoulou [41] achieved it for *-subalgebras $\mathcal{A}$ of $\mathrm{M}_{n}$, and we do it for real *-subalgebras $\mathcal{A}$ of $\mathrm{M}_{n}$ in Sec. 5 using Clausing’s work on retractions [16]. The analogue of Thm. 1.2 is established for the algebra $\mathcal{A}=\mathrm{M}_{n}(\mathbb{R})$ of real $n\times n$ matrices in Sec. 7. The Secs. 2 and 3 collect preliminaries.
Remark 1.3.
  1. a)
    Vesterstrøm [51] proved that the restriction of states to the center of a von Neumann algebra is open. Thm. 1.2 is a noncommutative analogue in the finite-dimensional setting.
  2. b)
    Stability is a meaningful concept in optimal control [41, 57] and quantum information theory [48] because of the following “CE-property”. A compact convex set $K$ is stable if and only if for every continuous function $f:K\to\mathbb{R}$, the envelope $f^{\vee}(x):=\sup\{h(x):h\leq f\}$, $x\in K$ is continuous, see [51, 39] and [48]. Here, the supremum is taken over all continuous affine functions $h:K\to\mathbb{R}$ whose graphs lie below the graph of $f$.
2. States and density matrices
The aim of this section is to translate openness questions from states to density matrices. In the sequel, we refer to matrix systems and *-subalgebras synonymously as complex matrix systems and complex *-subalgebras, respectively. We introduce their real counterparts because real *-subalgebras have a greater variety of state spaces than the complex ones.
A real matrix system on $\mathbb{C}^{n}$ is a real linear subspace $\mathcal{R}$ of $\mathrm{M}_{n}$ that is self-adjoint and contains $\mathds{1}_{n}$. Let $\mathcal{R}$ denote a real matrix system on $\mathbb{C}^{n}$. We endow $\mathcal{R}$ with the Euclidean scalar product
×$\mathcal{R}\times\mathcal{R}\to\mathbb{R}\,,\quad(A,B)\mapsto\mathrm{Re}\langle A,B\rangle\,,$ (2.1)
where $\langle A,B\rangle=\operatorname{tr}A^{\ast}B$ and $\mathrm{Re}(a+\mathrm{i}b)=a$ is the real part of a complex number, $A,B\in\mathrm{M}_{n}$, $a,b\in\mathbb{R}$. The positive cone, space of hermitian matrices, dual cone, and space of density matrices are defined verbatim to their respective complex counterparts defined in Sec. 1, and are denoted by
$\mathcal{C}(\mathcal{R}),\quad\mathcal{H}(\mathcal{R}),\quad\mathcal{C}(\mathcal{R})^{\vee},\quad\text{and}\quad\mathcal{D}(\mathcal{R})\,.$
We call $\mathcal{R}$ a real *-subalgebra of $\mathrm{M}_{n}$ if $\mathcal{R}$ is closed under matrix multiplication.
Generalizing a definition from real algebras [32, Sec. 4.5], we define the real state space of $\mathcal{R}$ as
$\mathcal{S}_{\mathbb{R}}(\mathcal{R}):=\{\ell\in\mathcal{R}_{\mathbb{R},0}^{\ast}\mid\forall A\in\mathcal{C}(\mathcal{R}):\ell(A)\geq 0,\ell(\mathds{1}_{n})=1\}\,,$
where
$\mathcal{R}_{\mathbb{R},0}^{\ast}:=\{\ell:\mathcal{R}\to\mathbb{R}\mid\mbox{$\ell$ is $\mathbb{R}$-linear},\forall A\in\mathcal{H}^{-}(\mathcal{R}):\ell(A)=0\}$
is the space of real functionals vanishing on the skew-hermitian matrices
():=$\mathcal{H}^{-}(\mathcal{R}):=\{A\in\mathcal{R}:A^{\ast}=-A\}\,.$
Lemma 2.1.
If $\mathcal{R}$ is a real matrix system on $\mathbb{C}^{n}$, then the map
$r_{\mathbb{R},\mathcal{R}}:\mathcal{D}(\mathcal{R})\to\mathcal{S}_{\mathbb{R}}(\mathcal{R}),\quad\rho\mapsto\mathrm{Re}\langle\rho,\,\cdot\,\rangle$
is a real affine isomorphism between compact convex sets.
Proof.
As $\mathcal{R}$ is the orthogonal direct sum $\mathcal{R}=\mathcal{H}(\mathcal{R})\oplus\mathcal{H}^{-}(\mathcal{R})$, the real linear isomorphism [23, Sec. 67]
$r_{\mathbb{R},\mathcal{R}}:\mathcal{R}\to\{\ell:\mathcal{R}\to\mathbb{R}\mid\mbox{$\ell$ is $\mathbb{R}$-linear}\}\,,\quad A\mapsto\mathrm{Re}\langle A,\,\cdot\,\rangle$
restricts to the real linear isomorphism $\mathcal{H}(\mathcal{R})\to\mathcal{R}_{\mathbb{R},0}^{\ast}$. Restricting $r_{\mathbb{R},\mathcal{R}}$ further, we obtain an injective map whose domain is $\mathcal{D}(\mathcal{R})$. So, it suffices to prove $r_{\mathbb{R},\mathcal{R}}(\rho)\in\mathcal{S}_{\mathbb{R}}(\mathcal{R})$ for $\rho\in\mathcal{D}(\mathcal{R})$, and that $r_{\mathbb{R},\mathcal{R}}:\mathcal{D}(\mathcal{R})\to\mathcal{S}_{\mathbb{R}}(\mathcal{R})$ is surjective. Both assertions are straightforward to verify. The convex set $\mathcal{D}(\mathcal{R})$ is compact, see [44, Lemma 3.3], because $\mathds{1}_{n}$ lies in the interior of $\mathcal{C}(\mathcal{R})$ in the topology of $\mathcal{H}(\mathcal{R})$. The convex set $\mathcal{S}_{\mathbb{R}}(\mathcal{R})$ is compact as it is the image of a compact set under a continuous map. ∎
Returning to complex functionals, we consider the real vector space
$\mathcal{R}_{\mathrm{her}}^{\ast}:=\{\ell:\mathcal{R}\to\mathbb{C}\mid\mbox{$\ell$ is $\mathbb{C}$-linear},\forall A\in\mathcal{H}(\mathcal{R}):\ell(A)\in\mathbb{R}\}$
of complex functionals taking real values on hermitian matrices.
Lemma 2.2.
If $\mathcal{R}$ is a complex matrix system on $\mathbb{C}^{n}$, then the map
$r_{\mathcal{R}}:\mathcal{D}(\mathcal{R})\to\mathcal{S}(\mathcal{R})\,,\quad\rho\mapsto\langle\rho,\,\cdot\,\rangle$
is a real affine isomorphism between compact convex sets.
Proof.
The complex antilinear isomorphism $r_{\mathcal{R}}:\mathcal{R}\to\mathcal{R}^{\ast}$, $A\mapsto\langle A,\,\cdot\,\rangle$ restricts to a real linear isomorphism $\mathcal{H}(\mathcal{R})\to\mathcal{R}_{\mathrm{her}}^{\ast}$. Furthermore,
$\alpha:\mathcal{R}_{\mathbb{R},0}^{\ast}\to\mathcal{R}_{\mathrm{her}}^{\ast}\,,\quad\alpha(\ell)[A+\mathrm{i}B]:=\ell(A)+\mathrm{i}\ell(B)\,,\quad A,B\in\mathcal{H}(\mathcal{R})\,,$
is a real linear isomorphism, whose inverse is given by $\ell\mapsto\mathrm{Re}\circ\ell$. So, the following diagram commutes. (Note that two arrows in opposite directions denote a bijection.)
${\mathcal{H}(\mathcal{R})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{r_{\mathcal{R}}}$$\scriptstyle{r_{\mathbb{R},\mathcal{R}}}$${\mathcal{R}_{\mathbb{R},0}^{\ast}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\alpha}$${\mathcal{R}_{\mathrm{her}}^{\ast}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathrm{Re}}$
Moreover, if $\ell_{1}\in\mathcal{R}_{\mathbb{R},0}^{\ast}$ and $\ell_{2}\in\mathcal{R}_{\mathrm{her}}^{\ast}$ satisfy $\ell_{2}=\alpha(\ell_{1})$, then $\ell_{1}\in\mathcal{S}_{\mathbb{R}}(\mathcal{R})$ holds if and only if $\ell_{2}\in\mathcal{S}(\mathcal{R})$. Hence, the claim follows from Lemma 2.1. ∎
Example 2.3.
Real *-subalgebras of $\mathrm{M}_{n}$ have a richer class of state spaces than the complex ones. The Bloch ball [7, Sec. 5.2]
$\mathcal{D}(\mathrm{M}_{2})=\big{\{}\frac{1}{2}(\mathds{1}_{2}+c_{X}X+c_{Y}Y+c_{Z}Z)\colon c_{X},c_{Y},c_{Z}\in\mathbb{R},c_{X}^{2}+c_{Y}^{2}+c_{Z}^{2}\leq 1\big{\}}$
is a three-dimensional Euclidean ball of radius $1/\sqrt{2}$. The set of density matrices of the algebra $\mathrm{M}_{2}(\mathbb{R})$ of real $2\times 2$ matrices is the great disk
$\mathcal{D}(\mathrm{M}_{2}(\mathbb{R}))=\big{\{}\frac{1}{2}(\mathds{1}_{2}+c_{X}X+c_{Z}Z)\colon c_{X},c_{Z}\in\mathbb{R},c_{X}^{2}+c_{Z}^{2}\leq 1\big{\}}$
of the Bloch ball $\mathcal{D}(\mathrm{M}_{2})$. There is no complex *-subalgebra of $\mathrm{M}_{n}$ in any dimension $n$ whose state space is a disk.
Remark 2.4.
Real and complex matrix systems have the same families of state spaces. If $\mathcal{R}$ is a real or complex matrix system on $\mathbb{C}^{n}$, then the real matrix system of hermitian matrices $\mathcal{H}(\mathcal{R})$ has the same set of hermitian matrices and hence the same set of density matrices as $\mathcal{R}$. Conversely, any real matrix system $\mathcal{R}$ included in $\mathcal{H}(\mathrm{M}_{n})$ has the same same set of hermitian matrices and the same set of density matrices as the complex matrix system $\mathcal{R}\oplus\mathrm{i}\mathcal{R}$. The set of density matrices of a real or complex matrix system $\mathcal{R}$ is affinely isomorphic the state space of $\mathcal{R}$ by Lemma 2.1 or 2.2, respectively.
Let $\mathcal{R}_{1},\mathcal{R}_{2}$ be real matrix systems on $\mathbb{C}^{n}$ and let $\mathcal{R}_{2}\subset\mathcal{R}_{1}$. We abbreviate $\mathcal{H}_{i}:=\mathcal{H}(\mathcal{R}_{i})$, $\mathcal{C}_{i}:=\mathcal{C}(\mathcal{R}_{i})$, $\mathcal{C}_{i}^{\vee}:=\mathcal{C}^{\vee}(\mathcal{R}_{i})$, $\mathcal{D}_{i}:=\mathcal{D}(\mathcal{R}_{i})$, $\mathcal{S}_{\mathbb{R},i}:=\mathcal{S}_{\mathbb{R}}(\mathcal{R}_{i})$, $\mathcal{S}_{i}:=\mathcal{S}(\mathcal{R}_{i})$, $r_{\mathbb{R},i}:=r_{\mathbb{R},\mathcal{R}_{i}}$, and $r_{i}:=r_{\mathcal{R}_{i}}$, $i=1,2$. Usually, the orthogonal projection of $\mathcal{R}_{1}$ onto $\mathcal{R}_{2}$ is the idempotent self-adjoint linear map $\mathcal{R}_{1}\to\mathcal{R}_{1}$ whose range is $\mathcal{R}_{2}$. Reducing the codomain to the range, we get a map
$\pi:\mathcal{R}_{1}\to\mathcal{R}_{2}\,,$ (2.2)
whose value at $A\in\mathcal{R}_{1}$ is specified by the equations $\mathrm{Re}\langle A-\pi(A),B\rangle=0$ for all $B\in\mathcal{R}_{2}$. We refer to $\pi$ as the orthogonal projection of $\mathcal{R}_{1}$ onto $\mathcal{R}_{2}$ in this paper. The notation $\mathcal{R}_{1}\to\mathcal{R}_{2}$ conveying domain and range is useful especially in Sec. 6. The adjoint of $\pi$ is the embedding $\mathcal{R}_{2}\to\mathcal{R}_{1}$, $A\mapsto A$. If $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$ are complex matrix systems, then the Frobenius inner product induces the same orthogonal projection as the Euclidean scalar product (2.1).
As $\mathcal{R}_{i}$ is the orthogonal direct sum of the spaces of its hermitian and skew-hermitian matrices, $i=1,2$, the map (2.2) restricts to $\pi:\mathcal{H}_{1}\to\mathcal{H}_{2}$, which we denote (aware of the notational imprecision) by the same symbol $\pi$. The value of $\pi$ at $A\in\mathcal{H}_{1}$ is specified by $\langle A-\pi(A),B\rangle=0$ for all $B\in\mathcal{H}_{2}$.
A cone in a Euclidean space $(E,\langle\!\langle\cdot,\cdot\rangle\!\rangle)$ is a subset $C$ of $E$ that is closed under multiplication with positive scalars. A base of a cone $C$ is a subset $B$ of $C$ such that $0\not\in\operatorname{aff}B$ and such that for all nonzero $x\in C$ there exist $y\in B$ and $s>0$ such that $x=sy$ holds. Note that we have $B=C\cap\operatorname{aff}B$ for every base $B$ of a cone $C$. The set $M^{\vee}:=\{x\in E\mid\forall y\in M\colon\langle\!\langle x,y\rangle\!\rangle\geq 0\}$ is a closed convex cone for every subset $M\subset E$, called the dual cone to $M$. Regarding duality of convex cones, we refer to [45, Sec. 14].
Lemma 2.5.
Let $\mathcal{R}_{1},\mathcal{R}_{2}$ be real matrix systems on $\mathbb{C}^{n}$ such that $\mathcal{R}_{2}\subset\mathcal{R}_{1}$ and let $\pi:\mathcal{H}_{1}\to\mathcal{H}_{2}$ denote the orthogonal projection. Then $\mathcal{C}_{2}^{\vee}=\pi(\mathcal{C}_{1}^{\vee})$ and $\mathcal{D}_{2}=\pi(\mathcal{D}_{1})$ holds.
Proof.
This lemma and its proof are similar to [44, Prop. 5.2]. The inclusions $\pi(\mathcal{C}_{1}^{\vee})\subset\mathcal{C}_{2}^{\vee}$ and $\mathcal{C}_{2}\supset(\pi(\mathcal{C}_{1}^{\vee}))^{\vee}$ are straightforward to verify and imply
$\pi(\mathcal{C}_{1}^{\vee})\subset\mathcal{C}_{2}^{\vee}\subset[\pi(\mathcal{C}_{1}^{\vee})]^{\vee}{}^{\vee}\,.$
Since $[\pi(\mathcal{C}_{1}^{\vee})]^{\vee}{}^{\vee}$ is the closure of the convex cone $\pi(\mathcal{C}_{1}^{\vee})$, it suffices to show that $\pi(\mathcal{C}_{1}^{\vee})$ is closed. By [44, Lemma 3.1], this would follow if $\pi(\mathcal{D}_{1})$ was a compact base of $\pi(\mathcal{C}_{1}^{\vee})$. As $\mathds{1}_{n}$ lies in the interior of $\mathcal{C}_{1}$ in the topology of $\mathcal{H}_{1}$, we know that $\mathcal{D}_{1}$ is a compact base of $\mathcal{C}_{1}^{\vee}$, see [44, Lemma 3.3]. Hence $\pi(\mathcal{D}_{1})$ is a compact base of $\pi(\mathcal{C}_{1}^{\vee})$ provided we establish that $0\not\in\operatorname{aff}\pi(\mathcal{D}_{1})$. But this is clear from $\pi$ being trace-preserving:
$\operatorname{tr}(\pi(A))=\langle\mathds{1}_{n},\pi(A)\rangle=\langle\mathds{1}_{n},A\rangle=\operatorname{tr}(A)\,,\quad A\in\mathcal{H}_{1}\,.$
This completes the proof of $\mathcal{C}_{2}^{\vee}=\pi(\mathcal{C}_{1}^{\vee})$. The identity $\mathcal{C}_{2}^{\vee}=\pi(\mathcal{C}_{1}^{\vee})$ and the fact that $\pi$ is trace-preserving imply $\mathcal{D}_{2}=\pi(\mathcal{D}_{1})$. ∎
Lemma 2.5 shows that the orthogonal projection $\pi:\mathcal{R}_{1}\to\mathcal{R}_{2}$ restricts to the map
$\pi:\mathcal{D}_{1}\to\mathcal{D}_{2}\,,$ (2.3)
which we call the orthogonal projection of $\mathcal{D}_{1}$ onto $\mathcal{D}_{2}$. To avoid any possible confusion, the orthogonal projections $\mathcal{R}_{1}\to\mathcal{R}_{2}$ and $\mathcal{H}_{1}\to\mathcal{H}_{2}$ are written without function labels from here on (Lemma 3.7 is an exception).
Proposition 2.6.
Let $\mathcal{R}_{1},\mathcal{R}_{2}$ be real matrix systems on $\mathbb{C}^{n}$ such that $\mathcal{R}_{2}\subset\mathcal{R}_{1}$ and let $\pi:\mathcal{D}(\mathcal{R}_{1})\to\mathcal{D}(\mathcal{R}_{2})$ denote the orthogonal projection.
  1. a)
    The diagram a) below commutes. Both $\pi:\mathcal{D}_{1}\to\mathcal{D}_{2}$ and $\mathcal{S}_{\mathbb{R},1}\to\mathcal{S}_{\mathbb{R},2}$, $\ell\mapsto\ell|_{\mathcal{R}_{2}}$ are surjective maps.
  2. b)
    If $\mathcal{R}_{1},\mathcal{R}_{2}$ are complex matrix systems on $\mathbb{C}^{n}$, then the diagram b) below commutes and the maps $\pi:\mathcal{D}_{1}\to\mathcal{D}_{2}$ and $\mathcal{S}_{1}\to\mathcal{S}_{2}$, $\ell\mapsto\ell|_{\mathcal{R}_{2}}$ are onto.
a)   ${\mathcal{D}_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{r_{\mathbb{R},1}}$$\scriptstyle{\pi}$${\mathcal{S}_{\mathbb{R},1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\ell\mapsto\ell|_{\mathcal{R}_{2}}}$${\mathcal{D}_{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{r_{\mathbb{R},2}}$${\mathcal{S}_{\mathbb{R},2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$            b)   ${\mathcal{D}_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{r_{1}}$$\scriptstyle{\pi}$${\mathcal{S}_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\ell\mapsto\ell|_{\mathcal{R}_{2}}}$${\mathcal{D}_{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{r_{2}}$${\mathcal{S}_{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
Proof.
a) The horizontal arrows of diagram a) and b) are obtained in Lemma 2.1 and Lemma 2.2, respectively, and $\pi:\mathcal{D}_{1}\to\mathcal{D}_{2}$ is onto by Lemma 2.5. The diagram a) commutes because for all $\rho\in\mathcal{D}_{1}$ and $A\in\mathcal{R}_{2}$ we have
$[r_{\mathbb{R},2}\circ\pi(\rho)](A)=\mathrm{Re}\langle\pi(\rho),A\rangle=\mathrm{Re}\langle\rho,A\rangle=[r_{\mathbb{R},1}(\rho)](A)=r_{\mathbb{R},1}(\rho)|_{\mathcal{R}_{2}}(A)\,.$
Therefore and since $\pi:\mathcal{D}_{1}\to\mathcal{D}_{2}$ is surjective, the map $\mathcal{S}_{\mathbb{R},1}\to\mathcal{S}_{\mathbb{R},2}$ is surjective as well. The proof of b) is similar. ∎
Corollary 2.7.
Let $\mathcal{R}_{1},\mathcal{R}_{2}$ be real matrix systems on $\mathbb{C}^{n}$ such that $\mathcal{R}_{2}\subset\mathcal{R}_{1}$, let $\pi:\mathcal{D}(\mathcal{R}_{1})\to\mathcal{D}(\mathcal{R}_{2})$ denote the orthogonal projection, and let $\rho\in\mathcal{D}_{1}$.
  1. a)
    The map $\pi:\mathcal{D}_{1}\to\mathcal{D}_{2}$ is open at $\rho$ if and only if $\mathcal{S}_{\mathbb{R},1}\to\mathcal{S}_{\mathbb{R},2}$, $\ell\mapsto\ell|_{\mathcal{R}_{2}}$ is open at $r_{\mathbb{R},1}(\rho)$.
  2. b)
    If $\mathcal{R}_{1},\mathcal{R}_{2}$ are complex matrix systems on $\mathbb{C}^{n}$, then $\pi:\mathcal{D}_{1}\to\mathcal{D}_{2}$ is open at $\rho$ if and only if $\mathcal{S}_{1}\to\mathcal{S}_{2}$, $\ell\mapsto\ell|_{\mathcal{R}_{2}}$ is open at $r_{1}(\rho)$.
Proof.
This follows directly from Prop. 2.6. ∎
Remark 2.8.
Let $\mathcal{R}_{1},\mathcal{R}_{2}$ be real matrix systems on $\mathbb{C}^{n}$ and let $\mathcal{R}_{2}\subset\mathcal{R}_{1}$. Then $\mathcal{D}_{2}\supset\mathcal{D}_{1}\cap\mathcal{R}_{2}$ holds, but the converse inclusion is wrong in general, as Ex. 2.9 shows. However, if $\mathcal{R}_{2}$ is a real *-subalgebra of $\mathrm{M}_{n}$, then we have
$\mathcal{D}_{2}=\mathcal{D}_{1}\cap\mathcal{R}_{2}\,.$ (2.4)
Indeed, $\mathcal{C}_{2}^{\vee}=\mathcal{C}_{2}$ holds [28, Thm. III.2.1] as the space of hermitian matrices $\mathcal{H}_{2}$ is a Euclidean Jordan algebra with Jordan product $A\circ B=\frac{1}{2}(AB+BA)$ and inner product $(A,B)\mapsto\mathrm{Re}\langle A,B\rangle$, $A,B\in\mathcal{H}_{2}$. So, $\mathcal{C}_{2}^{\vee}=\mathcal{C}_{2}\subset\mathcal{C}_{1}\subset\mathcal{C}_{1}^{\vee}$ proves $\mathcal{C}_{2}^{\vee}=\mathcal{C}_{2}$, which implies (2.4). See also the Notes to Chapter 6 in [2].
Example 2.9.
Despite $\mathcal{R}_{2}\subset\mathcal{R}_{1}\subset\mathrm{M}_{3}$, the inclusions $\mathcal{D}_{2}\subset\mathcal{D}_{1}\subset\mathcal{D}(\mathrm{M}_{3})$ fail if $\mathcal{R}_{2}:=\operatorname{span}_{\mathbb{C}}(\mathds{1}_{3},Z\oplus 0)$ and $\mathcal{R}_{1}:=\operatorname{span}_{\mathbb{C}}(\mathds{1}_{3},X\oplus 1,Z\oplus 0)$. We have
$A_{\lambda}\in\mathcal{D}_{2}\Leftrightarrow|\lambda|\leq\frac{3}{2}\,,\quad A_{\lambda}\in\mathcal{D}_{1}\Leftrightarrow|\lambda|\leq\sqrt{2}\,,\quad A_{\lambda}\in\mathcal{D}(\mathrm{M}_{3})\Leftrightarrow|\lambda|\leq 1$
for $A_{\lambda}:=(\mathds{1}_{3}+\lambda\,Z\oplus 0)/3\in\mathcal{R}_{2}$, $\lambda\in\mathbb{R}$. The second equivalence is obtained by minimizing $\langle A_{\lambda},A\rangle$ for fixed $\lambda$ over $A\in\mathcal{C}_{1}$, that is, by minimizing $\langle A_{\lambda},\mathds{1}_{3}+c_{1}(X\oplus 1)+c_{2}Z\oplus 0\rangle=1+\frac{1}{3}c_{1}+\frac{2}{3}\lambda c_{2}$ on the unit disk of $\mathbb{R}^{2}$.
3. Direct convex sums
This section addresses affinely independent convex sets, their convex hulls, and maps defined thereon. Let
$\Delta_{m}:=\{(s_{1},\dots,s_{m})\in\mathbb{R}^{m}\mid\forall i:s_{i}\geq 0,s_{1}+\dots+s_{m}=1\}$
denote the probability simplex, and
$\Delta_{m}(\epsilon,s_{1},\dots,s_{m}):=\Delta_{m}\cap\bigoplus_{i=1}^{m}(s_{i}-\epsilon,s_{i}+\epsilon)$
the open hypercube of edge length $2\epsilon$ centered at $(s_{1},\dots,s_{m})\in\Delta_{m}$.
A family of convex subsets $K_{1},\dots,K_{m}$ of a Euclidean space is affinely independent if every point in their convex hull can be expressed by a unique convex combination $s_{1}x_{1}+\dots+s_{m}x_{m}$. This means that $(s_{1},\dots,s_{m})\in\Delta_{m}$ is unique and $x_{i}\in K_{i}$ is unique for all $i$ for which if $s_{i}>0$, $i=1,\dots,m$. The direct convex sum [1] of a family $K_{1},\dots,K_{m}$ of affinely independent convex sets is defined as their convex hull
$K_{1}\oplus_{\mathrm{c}}\cdots\oplus_{\mathrm{c}}K_{m}:=\operatorname{conv}(K_{1}\cup\dots\cup K_{m})\,.$
If $K_{1},\dots,K_{m}$ is a family of affinely independent compact convex sets, then their direct convex sum is compact [45, Thm. 17.2]. A compactness argument allows us to describe a base of open neighborhoods.
Lemma 3.1.
Let $K_{1},\dots,K_{m}$ be affinely independent compact convex subsets of a Euclidean space, and let $x_{i}\in K_{i}$, $i=1,\dots,m$, and $(s_{1},\dots,s_{m})\in\Delta_{m}$. Let $I:=\{i\in\{1,\ldots,m\}\mid s_{i}>0\}$, $\delta:=\min_{i\in I}s_{i}$, and define
$O_{I}(\epsilon,(A_{i})_{i\in I}):=\{t_{1}y_{1}+\dots+t_{m}y_{m}\mid\hskip 2.0&(t_{1},\dots,t_{m})\in\Delta_{m}(\epsilon,s_{1},\dots,s_{m}),\\ \hskip 2.0&\mbox{$\forall i\colon y_{i}\in K_{i}$ and ($i\in I\Rightarrow y_{i}\in A_{i}$)}\}\end{array}$
for every $\epsilon\in(0,\delta]$ and open set $A_{i}$ in the relative topology of $K_{i}$ containing $x_{i}$, $i=1,\dots,m$. Then the family $\{O_{I}(\epsilon,(A_{i})_{i\in I})\}$ is a local base of open neighborhoods at $s_{1}x_{1}+\dots+s_{m}x_{m}$ in the relative topology of $K_{1}\oplus_{\mathrm{c}}\cdots\oplus_{\mathrm{c}}K_{m}$.
Proof.
As $K:=K_{1}\oplus_{\mathrm{c}}\cdots\oplus_{\mathrm{c}}K_{m}$ is a metric space, it suffices to show that there are arbitrary small family members and $O:=O_{I}(\epsilon,(A_{i})_{i\in I})$ is open.
Let $B_{i}:=A_{i}$ if $i\in I$ and $B_{i}:=K_{i}$ if $i\not\in I$, $i=1,\dots,m$. Then the set
$U:=\Delta_{m}(\epsilon,s_{1},\dots,s_{m})\oplus\bigoplus_{i}B_{i}$
is open in the relative topology of the compact set $\widetilde{K}:=\Delta_{m}\oplus\bigoplus_{i}K_{i}$. The complement
$U^{\complement}:=\widetilde{K}\setminus U=\underbrace{\textstyle\Delta_{m}(\epsilon,s_{1},\dots,s_{m})^{\complement}\oplus\bigoplus_{i}K_{i}}_{C:=}\cup\bigcup_{j\in I}\underbrace{\textstyle\Delta_{m}\oplus(\bigoplus_{i\neq j}K_{i})\oplus A_{j}^{\complement}}_{C_{j}:=}$
is compact. Hence, its image under the continuous surjective map
$\beta:\widetilde{K}\to K\,,\quad((t_{i})_{i=1}^{m},(y_{i})_{i=1}^{m})\mapsto\sum_{i=1}^{m}t_{i}y_{i}$
is compact. We prove that $O=\beta(U)$ is open in $K$ by showing that $\beta(U)$ is disjoint from $\beta(U^{\complement})$. Let
$u:=(t_{1},\dots,t_{m},y_{1},\dots,y_{m})\in U\,.$
Then $\beta(u)\not\in\beta(C)$ follows as the vector $(t_{i})_{i=1}^{m}$ is uniquely determined by $\beta(u)$. For every $j\in I$ we have $s_{j}>0$, hence $t_{j}>0$ by the definition of $O$. Thus, $y_{j}$ is uniquely determined by $\beta(u)$, which proves $\beta(u)\not\in\beta(C_{j})$.
Let $|M|:=\sup_{x,y\in M}|y-x|$ denote the diameter of a set $M$. The distance of $x:=s_{1}x_{1}+\dots+s_{m}x_{m}$ from a point $t_{1}y_{1}+\dots+t_{m}y_{m}$ in $O$ is bounded by
$\sum_{i}\left(|s_{i}-t_{i}||x_{i}|+t_{i}|x_{i}-y_{i}|\right)\leq\epsilon\sum_{i}|x_{i}|+\sum_{i\in I}t_{i}|A_{i}|+\epsilon\sum_{i\not\in I}|K_{i}|\,.$
Choosing open sets $(A_{i})_{i\in I}$ with diameters at most $\epsilon$, we obtain
$|O|\leq 2\sup_{y\in O}|x-y|\leq 2\epsilon\left(\sum_{i}|x_{i}|+1+\sum_{i\not\in I}|K_{i}|\right)\,,$
which completes the proof, as the compact sets $K_{i}$ have finite diameters and as $\epsilon$ can be chosen arbitrarily small. ∎
Let $E_{1},\dots,E_{m}$ be Euclidean spaces. We consider $E_{i}$ as a subspace of the direct sum $\bigoplus_{j=1}^{m}E_{j}$ via the embedding
$E_{i}\to\bigoplus_{j=1}^{m}E_{j}\,,\quad x\mapsto(\underbrace{0,\dots,0}_{\makebox[0.0pt]{\scriptsize$i-1$ zeros}},x,\underbrace{0,\dots,0}_{\makebox[0.0pt]{\scriptsize$m-i$ zeros}})\,,\quad i=1,\dots,m\,.$
Proposition 3.2.
Let $K_{i}\subset E_{i}$ and $L_{i}\subset F_{i}$ be compact convex subsets of Euclidean spaces $E_{i}$ and $F_{i}$, $i=1,\dots,m$, such that $K_{1},\dots,K_{m}$ are affinely independent in $\bigoplus_{i=1}^{m}E_{i}$ and $L_{1},\dots,L_{m}$ are affinely independent in $\bigoplus_{i=1}^{m}F_{i}$. If $f_{i}:K_{i}\to L_{i}$ is a map, $i=1,\dots,m$, then a map
$f_{1}\oplus_{\mathrm{c}}\cdots\oplus_{\mathrm{c}}f_{m}:K_{1}\oplus_{\mathrm{c}}\cdots\oplus_{\mathrm{c}}K_{m}\to L_{1}\oplus_{\mathrm{c}}\cdots\oplus_{\mathrm{c}}L_{m}$ (3.1)
is well defined by
$s_{1}x_{1}+\dots+s_{m}x_{m}\mapsto s_{1}f_{1}(x_{1})+\dots+s_{m}f_{m}(x_{m})\,.$
If $f_{i}$ is open and surjective for $i=1,\dots,m$, then $f_{1}\oplus_{\mathrm{c}}\cdots\oplus_{\mathrm{c}}f_{m}$ is open.
Proof.
The map is well defined as the sets $K_{1},\dots,K_{m}$ are affinely independent and the sets $L_{1},\dots,L_{m}$ are affinely independent. Regarding the openness of $f_{1}\oplus_{\mathrm{c}}\cdots\oplus_{\mathrm{c}}f_{m}$, it suffices to show that the images of the members of a base of the relative topology of $K_{1}\oplus_{\mathrm{c}}\cdots\oplus_{\mathrm{c}}K_{m}$ are open. Using the base of Lemma 3.1, we have
$f_{1}\oplus_{\mathrm{c}}\cdots\oplus_{\mathrm{c}}f_{m}[O_{I}(\epsilon,(A_{i})_{i\in I})]=O_{I}[\epsilon,(f(A_{i}))_{i\in I}]\,,$
which is open in $L_{1}\oplus_{\mathrm{c}}\cdots\oplus_{\mathrm{c}}L_{m}$. The required identity of $f_{i}(K_{i})=L_{i}$ for every $i\not\in I$ is a consequence of $f_{i}$ being surjective. ∎
We call the map (3.1) the direct convex sum of the maps $f_{1},\dots,f_{m}$. Next, we recall a sufficient condition for the affine independence of convex sets [17].
Remark 3.3.
Let $K_{i}$ be a convex subset of a Euclidean space $E_{i}$ such that $0\not\in\operatorname{aff}K_{i}$ holds for $i=1,\dots,m$. Then $K_{1},\dots,K_{m}$ are affinely independent in the direct sum $\bigoplus_{i=1}^{m}E_{i}$.
Returning to matrix systems, we consider a real matrix system $\mathcal{R}_{i}$ on $\mathbb{C}^{n_{i}}$, $i=1,\dots,m$. The direct sum $\mathcal{R}:=\bigoplus_{i=1}^{m}\mathcal{R}_{i}$ is a real matrix system on $\mathbb{C}^{n_{1}+\dots+n_{m}}$. The Frobenius inner product of $(A_{i})_{i=1}^{m},(B_{i})_{i=1}^{m}\in\mathcal{R}$ is $\langle(A_{i})_{i=1}^{m},(B_{i})_{i=1}^{m}\rangle=\sum_{i=1}^{m}\langle A_{i},B_{i}\rangle$.
Lemma 3.4.
Let $\mathcal{R}_{i}$ be a real matrix system on $\mathbb{C}^{n_{i}}$, $i=1,\ldots,m$. Then
$\mathcal{C}(\mathcal{R}_{1}\oplus\dots\oplus\mathcal{R}_{m})$ $=\mathcal{C}(\mathcal{R}_{1})\oplus\dots\oplus\mathcal{C}(\mathcal{R}_{m})\,,$
$\mathcal{C}^{\vee}(\mathcal{R}_{1}\oplus\dots\oplus\mathcal{R}_{m})$ $=\mathcal{C}^{\vee}(\mathcal{R}_{1})\oplus\dots\oplus\mathcal{C}^{\vee}(\mathcal{R}_{m})\,,$
$\text{and}\quad\mathcal{D}(\mathcal{R}_{1}\oplus\dots\oplus\mathcal{R}_{m})$ $=\mathcal{D}(\mathcal{R}_{1})\oplus_{\mathrm{c}}\cdots\oplus_{\mathrm{c}}\mathcal{D}(\mathcal{R}_{m})\,.$
Proof.
The first identity is clear. The second one follows by induction from $m=2$, a case that is easy to verify. By Rem. 3.3 and because $\mathcal{D}(\mathcal{R}_{i})$ is included in the hyperplane of trace-one matrices $i=1,\dots,m$, the sets $\mathcal{D}(\mathcal{R}_{1}),\dots,\mathcal{D}(\mathcal{R}_{m})$ are affinely independent in the direct sum $\bigoplus_{i=1}^{m}\mathcal{H}(\mathcal{R}_{i})$. The third identity follows from the second one by enforcing the trace to be one. ∎
Example 3.5.
Let $\mathcal{R}:=\operatorname{span}_{\mathbb{R}}(\mathds{1}_{3},X\oplus 1,Z\oplus 0)$.
  1. a)
    The set of density matrices $\mathcal{D}(\mathrm{M}_{2}\oplus\mathrm{M}_{1})$ of the *-subalgebra $\mathrm{M}_{2}\oplus\mathrm{M}_{1}$ of $\mathrm{M}_{3}$ is a symmetric cone that is the direct convex sum of the Bloch ball $\mathcal{D}(\mathrm{M}_{2})$ and the singleton $\mathcal{D}(\mathrm{M}_{1})=\{1\}$, see Lemma 3.4 and Ex. 2.3. The closed segment
    $\mathcal{G}:=\big{[}\ket{+}\!\!\bra{+}\oplus 0,0\oplus 1\big{]}=\big{\{}(1-\lambda)\ket{+}\!\!\bra{+}\oplus\lambda\colon\lambda\in[0,1]\big{\}}$
    is a generatrix of this cone. One proves along the lines of Ex. 1.1 that the orthogonal projection
    $\pi:\mathcal{D}(\mathrm{M}_{2}\oplus\mathrm{M}_{1})\to\mathcal{D}(\mathcal{R})$
    is not open at any point in the half-open segment
    $\mathcal{G}_{0}:=\big{(}\ket{+}\!\!\bra{+}\oplus 0,0\oplus 1\big{]}=\mathcal{G}\setminus\{\ket{+}\!\!\bra{+}\oplus 0\}\,.$
    Thereby, the equivalence of states and density matrices is described in Coro. 2.7. It is important to observe that $\frac{1}{2}(\mathds{1}_{2}+\cos(\theta)X+\sin(\theta)Z)\oplus 0$ and $\cos(\theta)(X\oplus 1)+\sin(\theta)Z\oplus 0$ are matrices in $\mathrm{M}_{2}\oplus\mathrm{M}_{1}$ for all $\theta\in\mathbb{R}$. The map $\pi$ is open at every point in the complement of $\mathcal{G}_{0}$ because $\mathcal{D}(\mathrm{M}_{2}\oplus\mathrm{M}_{1})$ is a cone over a ball, see [52, Lemma 4.17] for a detailed proof.
  2. b)
    The orthogonal projection $\pi:\mathcal{D}(\mathrm{M}_{2}\oplus\mathrm{M}_{1})\to\mathcal{D}(\mathcal{R})$ has an instructive geometry. The set of density matrices $\mathcal{D}(\mathcal{R})=\pi\big{(}\mathcal{D}(\mathrm{M}_{2})\oplus_{\mathrm{c}}\{1\}\big{)}$ is the convex hull of the projected ball $\pi(\mathcal{D}(\mathrm{M}_{2})\oplus\{0\})$ and the singleton $\pi(0\oplus 1)$ by Lemma 2.5. In turn, $\pi(\mathcal{D}(\mathrm{M}_{2})\oplus\{0\})$ is the filled ellipse of all points
[…]
    $\textstyle=\frac{1}{2}\big{(}\mathds{1}_{3}-M+c_{X}(3M-\mathds{1}_{3})+c_{Z}Z\oplus 0\big{)}\,,$
    where $c_{X},c_{Y},c_{Z}\in\mathbb{R}$ satisfy $c_{X}^{2}+c_{Y}^{2}+c_{Z}^{2}=1$, and $M:=\frac{1}{2}(\ket{+}\!\!\bra{+}\oplus 1)$ is the midpoint of the generatrix $\mathcal{G}$. Note that $\mathds{1}_{3}$, $(3X-\mathds{1}_{2})\oplus 2$, $Z\oplus 0$ is an orthogonal basis of $\mathcal{R}$. The choice of $c_{X}=1,c_{Y}=c_{Z}=0$ yields
    $\pi(\ket{+}\!\!\bra{+}\oplus 0)=M=\pi(0\oplus 1)\,.$ (3.3)
    Thus, $\mathcal{G}$ is perpendicular to $\mathcal{R}$, and $\mathcal{D}(\mathcal{R})=\pi(\mathcal{D}(\mathrm{M}_{2})\oplus\{0\})$ is an ellipse. Moreover, $\pi^{-1}(M)=\mathcal{G}$ holds by equation (3.2) and (3.3).
  3. c)
    The ellipse $\mathcal{D}(\mathcal{R})$ has the semiaxes $\sqrt{3/8}$ and $1/\sqrt{2}$. This follows from the formula (3.2) when $(c_{X},c_{Y},c_{Z})$ is assigned the values of (±$(\pm 1,0,0)$ and $(0,0,\pm 1)$. In retrospect to Rem. 2.9, the value $(1/3,0,\sqrt{8/9})$ confirms that $(\mathds{1}_{3}+\lambda\,Z\oplus 0)/3$ is contained in $\mathcal{D}(\mathcal{R})$ if and only if $|\lambda|\leq\sqrt{2}$.
Example 3.6.
The set of density matrices $\mathcal{D}(\mathrm{M}_{2}(\mathbb{R})\oplus\mathrm{M}_{1}(\mathbb{R}))$ of the real *-subalgebra $\mathrm{M}_{2}(\mathbb{R})\oplus\mathrm{M}_{1}(\mathbb{R})$ of $\mathrm{M}_{3}$ is a symmetric cone, which is the direct convex sum of the great disk $\mathcal{D}(\mathrm{M}_{2}(\mathbb{R}))$ of the Bloch ball and a singleton. As in Ex. 3.5, the closed segment $\mathcal{G}$ is a generatrix of this cone, which is the fiber of the orthogonal projection
$\pi:\mathcal{D}(\mathrm{M}_{2}(\mathbb{R})\oplus\mathrm{M}_{1}(\mathbb{R}))\to\mathcal{D}(\mathcal{R})$
over $M$. The map $\pi$ is not open at any point in the half-open segment $\mathcal{G}_{0}$ and open at every point in the complement. The lack of openness can be visualized graphically in three-space by observing that $\pi$ projects the cone $\mathcal{D}(\mathrm{M}_{2}(\mathbb{R})\oplus\mathrm{M}_{1}(\mathbb{R}))$ along its generatrix $\mathcal{G}$ to the ellipse $\mathcal{D}(\mathcal{R})$.
Lemma 3.7.
Let $\mathcal{R}_{i},\mathcal{R}_{i}^{\prime}$ be real matrix systems on $\mathbb{C}^{n_{i}}$ such that $\mathcal{R}_{i}\subset\mathcal{R}_{i}^{\prime}$, let $\pi_{i}:\mathcal{R}_{i}^{\prime}\to\mathcal{R}_{i}$ denote the orthogonal projection, $i=1,\ldots,m$, and consider the direct sums $\mathcal{R}^{\prime}:=\bigoplus_{i=1}^{m}\mathcal{R}_{i}^{\prime}$ and $\mathcal{R}:=\bigoplus_{i=1}^{m}\mathcal{R}_{i}$. The orthogonal projection $\pi:\mathcal{D}(\mathcal{R}^{\prime})\to\mathcal{D}(\mathcal{R})$ equals
$\pi_{1}\oplus_{\mathrm{c}}\cdots\oplus_{\mathrm{c}}\pi_{m}:\mathcal{D}(\mathcal{R}_{1}^{\prime})\oplus_{\mathrm{c}}\cdots\oplus_{\mathrm{c}}\mathcal{D}(\mathcal{R}_{m}^{\prime})\to\mathcal{D}(\mathcal{R}_{1})\oplus_{\mathrm{c}}\cdots\oplus_{\mathrm{c}}\mathcal{D}(\mathcal{R}_{m})\,.$
Proof.
The orthogonal projections $\pi_{i}:\mathcal{D}(\mathcal{R}_{i}^{\prime})\to\mathcal{D}(\mathcal{R}_{i})$, $i=1,\ldots,m$, and $\pi:\mathcal{D}(\mathcal{R}^{\prime})\to\mathcal{D}(\mathcal{R})$ are well defined by Lemma 2.5. A straight-forward computation shows that the orthogonal projection $\pi:\mathcal{R}^{\prime}\to\mathcal{R}$ is the direct sum $\bigoplus_{i=1}^{m}\pi_{i}$. The claim then follows from the third identity of Lemma 3.4 and from the definition of the direct convex sum of maps in formula (3.1). ∎
4. Numerical ranges
The orthogonal projection $\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathcal{R})$ may be restricted to the set $\operatorname{ex}\mathcal{D}(\mathrm{M}_{n})$ of extreme points [^1] An point in a convex set $K$ is an extreme point [45] of $K$ if there is no way to express it as a convex combination $(1-s)x+sy$ such that $x,y\in K$ and $s\in(0,1)$, except by taking $x=y$. We denote the set of extreme points of $K$ by $\operatorname{ex}K$. of $\mathcal{D}(\mathrm{M}_{n})$. The openness of this restriction was studied in matrix theory [18, 29, 30, 31, 34, 35, 49, 50, 53] for two-dimensional state spaces $\mathcal{D}(\mathcal{R})$ represented as numerical ranges (see Rem. 4.3).
Let $A_{1},\ldots,A_{k}\in\mathcal{H}(\mathrm{M}_{n})$ and consider the real matrix system
$\mathcal{R}(A_{1},\ldots,A_{k}):=\operatorname{span}_{\mathbb{R}}(\mathds{1}_{n},A_{1},\ldots,A_{k})\,.$
The image of $\mathcal{D}(\mathrm{M}_{n})$ under the real linear map
$v:\mathcal{H}(\mathrm{M}_{n})\to\mathbb{R}^{k}\,,\quad B\mapsto(\langle B,A_{1}\rangle,\dots,\langle B,A_{k}\rangle){}^{\mathrm{T}}$
is the joint numerical range $V(A_{1},\ldots,A_{k}):=v(\mathcal{D}(\mathrm{M}_{n}))\subset\mathbb{R}^{k}$, see [9]. Aware of the notational imprecision, we use the same label $v$ also for several restrictions of $v$, among others for
$v:\mathcal{D}(\mathrm{M}_{n})$ $\to V(A_{1},\ldots,A_{k})\,,$ $\rho$ $\mapsto(\langle\rho,A_{1}\rangle,\dots,\langle\rho,A_{k}\rangle){}^{\mathrm{T}}\,.$ (4.1)
Lemma 4.1.
The following diagram commutes.
${\mathcal{D}(\mathrm{M}_{n})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi}$$\scriptstyle{v}$${V(A_{1},\ldots,A_{k})}$${\mathcal{D}(\mathcal{R}(A_{1},\ldots,A_{k}))}$$\scriptstyle{v}$
Proof.
Let $\mathcal{R}:=\mathcal{R}(A_{1},\ldots,A_{k})$. By Lemma 2.5, the orthogonal projection $\pi:\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathcal{R})$ is surjective. It is straightforward to verify that $v:\mathcal{H}(\mathrm{M}_{n})\to\mathbb{R}^{k}$ factors through $\mathcal{R}$, in the sense that $v=v\circ\pi$ holds. Hence, the commutativity of the diagram is implied by the injectivity of $v$ restricted to the affine space $\{B\in\mathcal{H}(\mathcal{R}):\operatorname{tr}(B)=1\}$. Let $B_{1},B_{2}$ be contained in this affine space. If $v(B_{1})=v(B_{2})$, then
00 0 $=v(B_{1}-B_{2})=\langle B_{1}-B_{2},A_{i}\rangle_{i=1}^{k}$
[…]
This implies $B_{1}-B_{2}=0$ as $B_{1}-B_{2}\in\mathcal{H}(\mathcal{R})=\operatorname{span}_{\mathbb{R}}(\mathds{1}_{n},A_{1},\ldots,A_{k})$. ∎
In the remainder of this section, let $k=2$ and $A:=A_{1}+\mathrm{i}A_{2}$. The image of the unit sphere $\mathbb{C}S^{n}:=\{\ket{\varphi}\in\mathbb{C}^{n}:\braket{\varphi}{\varphi}=1\}$ under the hermitian quadratic form $f_{A}:\mathbb{C}^{n}\to\mathbb{C}$, |$\ket{\varphi}\mapsto\braket{\varphi}{A\varphi}$ is the numerical range
$W(A):=f_{A}(\mathbb{C}S^{n})\subset\mathbb{C}\,.$
Here, $\braket{\varphi_{1}}{\varphi_{2}}:=\overline{x_{1}}y_{1}+\cdots+\overline{x_{n}}y_{n}$ is the inner product of |$\ket{\varphi_{1}}=(x_{1},\ldots,x_{n}){}^{\mathrm{T}}$ and |$\ket{\varphi_{2}}=(y_{1},\ldots,y_{n}){}^{\mathrm{T}}$ in $\mathbb{C}^{n}$. We use the same label $f_{A}$ to denote the map
$f_{A}:\mathbb{C}S^{n}\to W(A)\,,\quad\ket{\varphi}\mapsto\braket{\varphi}{A\varphi}\,.$ (4.2)
Minkowski’s theorem [47] asserts that every compact convex set is the convex hull of its extreme points.
Proposition 4.2.
The following diagram commutes.
${\mathbb{C}S^{n}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$|$\scriptstyle{\ket{\varphi}\mapsto\ket{\varphi}\!\bra{\varphi}}$ |$\scriptstyle{f_{A}}$${\operatorname{ex}\mathcal{D}(\mathrm{M}_{n})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{v}$$\scriptstyle{\rho\mapsto\rho}$${\mathcal{D}(\mathrm{M}_{n})}$$\scriptstyle{v}$$\scriptstyle{\pi}$${W(A)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${V(A_{1},A_{2})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$z\mapsto\left(\begin{smallmatrix}\mathrm{Re}(z)\\ \mathrm{Im}(z)\end{smallmatrix}\right)$${\mathcal{D}(\mathcal{R}(A_{1},A_{2}))}$$\scriptstyle{v}$
Proof.
The bottom right triangle is the case $k=2$ of Lemma 4.1. The map $f_{A}:\mathbb{C}S^{n}\to W(A)$ factors through $\operatorname{ex}\mathcal{D}(\mathrm{M}_{n})$, $\mathcal{D}(\mathrm{M}_{n})$, and $V(A_{1},A_{2})$, as $\mathbb{C}S^{n}\to\mathcal{D}(\mathrm{M}_{n})$, |$\ket{\varphi}\mapsto\ket{\varphi}\!\!\bra{\varphi}$ | parametrizes the extreme points of $\mathcal{D}(\mathrm{M}_{n})$, see for example [7, Sec. 5.1], and since for all |$\ket{\varphi}\in\mathbb{C}S^{n}$ we have
[…]
It remains to show that $g:W(A)\to V(A_{1},A_{2})$, $z\mapsto(\mathrm{Re}(z),\mathrm{Im}(z)){}^{\mathrm{T}}$ is onto (being the restriction of a bijection, the map $g$ is one-to-one).
First, we show that $\operatorname{ex}V(A_{1},A_{2})$ is included in the image of $g$. The preimage of every extreme point $x$ of $V(A_{1},A_{2})$ under $v:\mathcal{D}(\mathrm{M}_{n})\to V(A_{1},A_{2})$ contains an extreme point of $\mathcal{D}(\mathrm{M}_{n})$. This is true because the preimage of $x$ is a face $\mathcal{F}$ of $\mathcal{D}(\mathrm{M}_{n})$, which has an extreme point $\rho$ by Minkowski’s theorem, since $\mathcal{D}(\mathrm{M}_{n})$, and hence $\mathcal{F}$, is compact. Since $\rho$ is also an extreme point of $\mathcal{D}(\mathrm{M}_{n})$, the claim follows from $\mathbb{C}S^{n}\to\operatorname{ex}\mathcal{D}(\mathrm{M}_{n})$ being onto.
Second, the convex hull of $\operatorname{ex}V(A_{1},A_{2})$ is included in the image of $g$ because $W(A)$ is convex by the Toeplitz-Hausdorff theorem [12, 37]. Third, the map $g$ is onto, because $V(A_{1},A_{2})$ is the convex hull of its extreme points, again by Minkowski’s theorem. ∎
The affine isomorphism $W(A)\cong\mathcal{D}(\mathcal{R}(A_{1},A_{2}))$ of Prop. 4.2 has an analogue in the much more general setting of matrix-valued states [22, Thm. 5.1].
Let $f_{A}^{-1}$ denote the multi-valued inverse of $f_{A}:\mathbb{C}S^{n}\to W(A)$. Corey et al. [18] define $f_{A}^{-1}$ to be strongly continuous at $z\in W(A)$ if the map $f_{A}$ is open at every point in the fiber $f_{A}^{-1}(z)$ of $f_{A}$ over $z$.
Remark 4.3.
Strong continuity can be described in terms of standard results on the numerical range. We refer to Sec. 8 of [29] and the references therein.
There exists a family of orthonormal bases |$\ket{\varphi_{1}(\theta)},\dots,\ket{\varphi_{n}(\theta)}$ of $\mathbb{C}^{n}$ that is analytically parametrized by a real number $\theta\in\mathbb{R}$; and there are analytic functions $\lambda_{i}:\mathbb{R}\to\mathbb{R}$, called eigenfunctions [30], such that
$(\cos(\theta)A_{1}+\sin(\theta)A_{2})\ket{\varphi_{i}(\theta)}=\lambda_{i}(\theta)\ket{\varphi_{i}(\theta)}\,,\quad\theta\in\mathbb{R}\,,\quad i=1,\dots,n\,.$
For every $i=1,\ldots,n$, the numerical range $W(A)$ includes the image $\operatorname{Img}(z_{i})$ of the curve
$z_{i}:\mathbb{R}\to\mathbb{C}\,,\quad\theta\mapsto e^{\mathrm{i}\theta}(\lambda_{i}(\theta)+\mathrm{i}\lambda^{\prime}_{i}(\theta))\,.$
Every extreme point of $W(A)$ is contained in $\operatorname{Img}(z_{i})$ for some $i=1,\ldots,n$.
If $z\in W(A)$ is not an extreme point of $W(A)$, then $f_{A}^{-1}$ is strongly continuous at $z$ [18, Thm. 4]. Let $z$ be an extreme point of $W(A)$. Then there are $\theta_{0}\in\mathbb{R}$ and $i_{0}\in\{1,\ldots,n\}$ such that $z=z_{i_{0}}(\theta_{0})$. Now, $f_{A}^{-1}$ is strongly continuous at $z$ if and only if $z_{i}(\theta_{0})=z$ implies $z_{i}=z_{i_{0}}$ for all $i=1,\ldots,n$, see [30, Thm. 2.1.1]. Leake at al. [30] state the latter condition by saying that the eigenfunctions corresponding to $z$ at $\theta_{0}$ do not split.
Remark 4.4.
Strong continuity is connected to the openness of a linear map. For all $z\in W(A)$, the multi-valued map $f_{A}^{-1}$ is strongly continuous at $z$ if and only if the restricted linear map $v:\mathcal{D}(\mathrm{M}_{n})\to V(A_{1},A_{2})$, introduced in (4.1), is open at every point in the fiber $v^{-1}[(\mathrm{Re}(z),\mathrm{Im}(z)){}^{\mathrm{T}}]$, see [53, Coro. 5.2]. Moreover, for every $x\in V(A_{1},A_{2})$, the map $v$ is open at some point in the relative interior [^2] The relative interior [45] of a convex set $K$, denoted by $\operatorname{ri}(K)$, is the interior of $K$ in the relative topology of the affine hull of $K$. of $v^{-1}(x)$ if and only if $v$ is open at every point in $v^{-1}(x)$.
Proposition 4.5.
Let $z\in W(A)$, let $x:=(\mathrm{Re}(z),\mathrm{Im}(z)){}^{\mathrm{T}}\in\mathbb{R}^{2}$, let $\rho$ be the unique density matrix of $\mathcal{R}(A_{1},A_{2})$ that satisfies $v(\rho)=x$, and let $\pi:\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathcal{R}(A_{1},A_{2}))$ denote the orthogonal projection. Then the following assertions are equivalent.
  • $f_{A}^{-1}$ is strongly continuous at $z$,
  • $v:\mathcal{D}(\mathrm{M}_{n})\to V(A_{1},A_{2})$ is open at every point in $v^{-1}(x)$,
  • $\pi:\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathcal{R}(A_{1},A_{2}))$ is open at every point in $\pi^{-1}(\rho)$.
We have $z\in\operatorname{ex}W(A)\Leftrightarrow x\in\operatorname{ex}V(A_{1},A_{2})\Leftrightarrow\rho\in\operatorname{ex}\mathcal{D}(\mathcal{R}(A_{1},A_{2}))$. If $z\in\operatorname{ex}W(A)$, then $z=z_{i_{0}}(\theta_{0})$ for some $\theta_{0}\in\mathbb{R}$ and $i_{0}\in\{1,\ldots,n\}$, and the following assertions are equivalent.
  • $f_{A}^{-1}$ is strongly continuous at $z$,
  • the eigenfunctions corresponding to $z$ at $\theta_{0}$ do not split,
  • $\pi$ is open at some point in the relative interior of $\pi^{-1}(\rho)$.
Proof.
The claims follow directly from Rem. 4.3 and 4.4, and Prop. 4.2. ∎
Example 4.6.
Prop. 4.5 ignores that $\pi:\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathcal{R}(A_{1},A_{2}))$ could be open at some point in the fiber $\pi^{-1}(\rho)$ over an extreme point $\rho$ of $\mathcal{D}(\mathcal{R}(A_{1},A_{2}))$, but not open anywhere in the relative interior of $\pi^{-1}(\rho)$.
This occurs for $A_{1}:=X\oplus 1$ and $A_{2}:=Z\oplus 0$. As discussed in Ex. 3.5 a) and b), the orthogonal projection $\pi_{\mathcal{R}}:\mathcal{D}(\mathrm{M}_{2}\oplus\mathrm{M}_{1})\to\mathcal{D}(\mathcal{R}(A_{1},A_{2}))$ is open at |++|$\ket{+}\!\!\bra{+}\oplus 0$ and nowhere else in the fiber $\pi_{\mathcal{R}}^{-1}(M)=\big{[}\ket{+}\!\!\bra{+}\oplus 0,0\oplus 1\big{]}$ over $M=\frac{1}{2}(\ket{+}\!\!\bra{+}\oplus 1)$. In particular, $\pi_{\mathcal{R}}$ is not open anywhere in the relative interior of $\pi_{\mathcal{R}}^{-1}(M)$. Ex. 8.3 proves the analogue for the larger fiber $\pi^{-1}(M)$.
5. Retractions
Generalizing a known result, we prove that the set of density matrices of every real *-subalgebra of $\mathrm{M}_{n}$ is a stable convex set. The proof relies on retractions of state spaces, a topic that also proved helpful in the foundations of quantum information theory [24].
Remark 5.1 (Stability of state spaces).
  1. a)
    The stability problem of a finite-dimensional compact convex set $K$ is completely solved [40]. The $d$-skeleton of $K$ is the union of all faces [^3] A face [45] of a convex set $K$ is a convex subset $F\subset K$ such that $x,y\in F$ is implied by $(1-s)x+sy\in F$ for all $x,y\in K$ and $s\in(0,1)$. of $K$ of dimension at most $d$. The convex set $K$ is stable if and only if for every nonnegative integer $d$, the $d$-skeleton of $K$ is closed.
  2. b)
    The set of density matrices $\mathcal{D}(\mathrm{M}_{n})$ is stable [20] because all its $d$-skeletons are closed. The closedness follows from three arguments: First, the set $\mathcal{D}(\mathrm{M}_{n})$ is a compact convex set of dimension $n^{2}-1$. Second, every nonempty face of $\mathrm{M}_{n}$ is unitarily similar to $\mathcal{D}(\mathrm{M}_{l})\oplus\{0\}$ for some positive integer $l$, and third, and the unitary group $U(n)$ is compact.
  3. c)
    The state space $\mathcal{S}(\mathcal{A})$ of every *-subalgebra $\mathcal{A}$ of $\mathrm{M}_{n}$ is stable [41] as it is a direct convex sum of state spaces of full matrix algebras $\mathrm{M}_{n}$. As $\mathcal{S}(\mathcal{A})$ is stable, the set of density matrices $\mathcal{D}(\mathcal{A})$ is stable, too, by Lemma 2.2.
A retraction is an affine map $f:K\to L$ between compact convex sets $K,L$ which is left-inverse to an affine map $g:L\to K$, called a section.
Proposition 5.2.
If $\mathcal{A}$ is a real *-subalgebra of $\mathrm{M}_{n}$, then the orthogonal projection $\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathcal{A})$ is a retraction and $\mathcal{D}(\mathcal{A})$ is stable.
Proof.
The orthogonal projection $\pi:\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathcal{A})$ is well defined by Lemma 2.5 and it is a retraction because the inclusion $\mathcal{D}(\mathcal{A})\subset\mathcal{D}(\mathrm{M}_{n})$ of equation (2.4) provides a section for $\pi$. Coro. 1.3 in [16] asserts that the image of a stable convex set under a retraction is stable. Therefore, and since $\mathcal{D}(\mathrm{M}_{n})$ is stable by Rem. 5.1 b), the convex set $\mathcal{D}(\mathcal{A})$ is stable. ∎
Example 5.3.
Prop. 5.2 does not generalize to arbitrary matrix systems. We consider the hermitian matrices
$A_{1}:=X\oplus 1\oplus 1\,,\quad A_{2}:=Z\oplus 0\oplus 0\,,\quad A_{3}:=0\oplus(-1)\oplus 1\,,$
in the algebra $\mathrm{M}_{2}\oplus\mathrm{M}_{1}\oplus\mathrm{M}_{1}$. The joint numerical range $V(A_{1},A_{2},A_{3})$, introduced in Sec. 4, is easier to handle algebraically than the set of density matrices $\mathcal{D}(\mathcal{R}(A_{1},A_{2},A_{3}))$, to which it is affinely isomorphic by Lemma 4.1. Since $A_{1},A_{2},A_{3}\in\mathrm{M}_{2}\oplus\mathrm{M}_{1}\oplus\mathrm{M}_{1}$, we have the standard result of
$V(A_{1},A_{2},A_{3})=\operatorname{conv}\big{(}V(X,Z,0)\cup V(1,0,-1)\cup V(1,0,1)\big{)}\,.$
Hence, $V(A_{1},A_{2},A_{3})=\operatorname{conv}(S)$ is the convex hull of
$S:=\{(c_{1},c_{2},0){}^{\mathrm{T}}\in\mathbb{R}^{3}\mid c_{1}^{2}+c_{2}^{2}=1\}\cup\{(1,0,-1){}^{\mathrm{T}},(1,0,1){}^{\mathrm{T}}\}\,.$
The set of extreme points $S\setminus\{(1,0,0){}^{\mathrm{T}}\}$ of $V(A_{1},A_{2},A_{3})$ is not closed, hence $V(A_{1},A_{2},A_{3})$ is not stable by Rem. 5.1 a).
6. Proof of Thm. 1.2
The main ideas in establishing Thm. 1.2 are that the set of density matrices $\mathcal{D}(\mathrm{M}_{n})$ is a stable convex set and that $\mathcal{D}(\mathrm{M}_{n})$ is highly symmetric. Being stable and symmetric, $\mathcal{D}(\mathrm{M}_{p+q})$ projects openly onto $\mathcal{D}(\mathrm{M}_{p}\oplus\mathrm{M}_{q})$. Loosely speaking, Thm. 1.2 is obtained by combining various such open projections, since every *-subalgebra of $\mathrm{M}_{n}$ is a direct sum of full matrix algebras.
If $K$ is a stable convex set, then the arithmetic mean map
$K^{\times k}\to K\,,\quad(x_{1},\dots,x_{k})\mapsto\frac{1}{k}\sum_{i=1}^{k}x_{i}$
the $k$-fold cartesian product of $K$ is open for all positive integers $k$. More generally, for every tuple $(s_{1},\dots,s_{k})$ of nonnegative real numbers adding up to one, the map
$K^{\times k}\to K\,,\quad(x_{1},\dots,x_{k})\mapsto\sum_{i=1}^{k}s_{i}x_{i}$
is open. The latter assertion is proved for $k=2$ in Prop. 1.1 in [17]. By induction, it is true for every $k>2$ as well.
Lemma 6.1.
Let $\mathcal{R}$ be a real matrix system on $\mathbb{C}^{n}$ and let $\mathcal{D}(\mathcal{R})$ be stable and invariant under an orthogonal transformation $\gamma:\mathcal{R}\to\mathcal{R}$ that generates a finite cyclic group. Let $\mathcal{A}:=\{A\in\mathcal{R}:\gamma(A)=A\}$ be a real *-subalgebra of $\mathrm{M}_{n}$. Then the orthogonal projection $\mathcal{D}(\mathcal{R})\to\mathcal{D}(\mathcal{A})$ is open.
Proof.
Let $k$ denote the order of the cyclic group generated by $\gamma$. The main idea is to write the orthogonal projection $\pi:\mathcal{D}(\mathcal{R})\to\mathcal{D}(\mathcal{A})$ in terms of the arithmetic mean map
$a:\mathcal{D}(\mathcal{R})^{\times k}\to\mathcal{D}(\mathcal{R})\,,\quad(\rho_{1},\dots,\rho_{k})\mapsto\frac{1}{k}\sum_{i=1}^{k}\rho_{i}\,,$
which, as discussed above, is open because $\mathcal{D}(\mathcal{R})$ is stable. The proof is done in two steps. First, we prove that
$a(\gamma(\mathcal{O})\times\dots\times\gamma^{k}(\mathcal{O}))\cap\mathcal{A}$ (6.1)
is an open subset of $\mathcal{D}(\mathcal{A})$ for all open subsets $\mathcal{O}$ of $\mathcal{D}(\mathcal{R})$. Secondly, we prove that
$a(\gamma(\mathcal{K})\times\dots\times\gamma^{k}(\mathcal{K}))\cap\mathcal{A}=\pi(\mathcal{K})$ (6.2)
holds for all convex subsets $\mathcal{K}$ of $\mathcal{D}(\mathcal{R})$. The assertions (6.1) and (6.2) together show that $\pi(\mathcal{O})$ is an open subset of $\mathcal{D}(\mathcal{A})$ for all convex open subsets $\mathcal{O}$ of $\mathcal{D}(\mathcal{R})$, and hence for all open subsets.
First, we prove that the set in (6.1) is open. As $\mathcal{D}(\mathcal{R})$ is invariant under the orthogonal transformation $\gamma$, the map $\gamma$ restricts to a homeomorphism $\mathcal{D}(\mathcal{R})\to\mathcal{D}(\mathcal{R})$. It follows that $\gamma(\mathcal{O})\times\dots\times\gamma^{k}(\mathcal{O})$ is an open subset of the $k$-fold cartesian product $\mathcal{D}(\mathcal{R})^{\times k}$. Then
$\widetilde{\mathcal{O}}:=a(\gamma(\mathcal{O})\times\dots\times\gamma^{k}(\mathcal{O}))$
is an open subset of $\mathcal{D}(\mathcal{R})$, because $\mathcal{D}(\mathcal{R})$ is stable. Finally, $\widetilde{\mathcal{O}}\cap\mathcal{A}$ is an open subset of $\mathcal{D}(\mathcal{A})$ by equation (2.4), which shows that $\widetilde{\mathcal{O}}\cap\mathcal{A}$ equals $\widetilde{\mathcal{O}}\cap\mathcal{D}(\mathcal{A})$.
Secondly, we prove the formula (6.2), beginning with the inclusion $\supset$”. Let $\rho\in\mathcal{K}\subset\mathcal{D}(\mathcal{R})$. The density matrix $\sigma:=a(\gamma(\rho),\dots,\gamma^{k}(\rho))$ is invariant under $\gamma$. This implies $\sigma\in\mathcal{A}$, hence $\sigma\in\mathcal{D}(\mathcal{A})$ by (2.4). Since $\gamma$ is self-adjoint, for all $A\in\mathcal{H}(\mathcal{A})$ we have
$\mathrm{Re}\langle\rho,A\rangle=\mathrm{Re}\langle a(\rho,\dots,\rho),A\rangle=\mathrm{Re}\langle a(\gamma(\rho),\dots,\gamma^{k}(\rho)),A\rangle=\mathrm{Re}\langle\sigma,A\rangle\,,$
hence $\pi(\rho)=\sigma$. Regarding the inclusion $\subset$”, let $\rho_{i}\in\mathcal{K}$, $i=1,\dots,k$. Let $\sigma:=a(\gamma(\rho_{1}),\dots,\gamma^{k}(\rho_{k}))$ and $\tau:=a(\rho_{1},\dots,\rho_{k})$. Then for all $A\in\mathcal{H}(\mathcal{A})$
$\mathrm{Re}\langle\sigma,A\rangle=\mathrm{Re}\langle a(\gamma(\rho_{1}),\dots,\gamma^{k}(\rho_{k})),A\rangle=\mathrm{Re}\langle a(\rho_{1},\dots,\rho_{k}),A\rangle=\mathrm{Re}\langle\tau,A\rangle$
holds. If $\sigma\in\mathcal{A}$, then $\sigma=\pi(\tau)$ follows. As $\mathcal{K}$ is convex, $\tau\in\mathcal{K}$ holds and we obtain $\sigma\in\pi(\mathcal{K})$. ∎
Remark 6.2.
A hermitian matrix $M\in\mathcal{H}(M_{p+q})$ in the block form
$M=\begin{pmatrix}A&B\\ B^{\ast}&C\end{pmatrix}\,,\quad A\in\mathcal{H}(M_{p})\,,\quad B\in\mathbb{C}^{p\times q}\,,\quad C\in\mathcal{H}(M_{q})$
is positive semidefinite if and only if the top left block $A$ is positive semidefinite, the range of $B$ is included in the range of $A$, and the generalized Schur complement $M/A=C-B^{\ast}A^{-}B$ is positive semidefinite, where $A^{-}$ is a generalized inverse of $A$, that is to say, $A^{-}\in M_{p}$ and $AA^{-}A=A$ holds [25].
Proposition 6.3.
The orthogonal projection $\mathcal{D}(\mathrm{M}_{p+q})\to\mathcal{D}(\mathrm{M}_{p}\oplus\mathrm{M}_{q})$ is open.
Proof.
Lemma 6.1 proves the claim when $\mathcal{R}:=\mathrm{M}_{p+q}$ and $\mathcal{A}:=\mathrm{M}_{p}\oplus\mathrm{M}_{q}$. The set of density matrices $\mathcal{D}(\mathrm{M}_{p+q})$ is stable by Rem. 5.1 b). The reflection $\gamma:\mathrm{M}_{p+q}\to\mathrm{M}_{p+q}$ at the subspace $\mathrm{M}_{p}\oplus\mathrm{M}_{q}$ generates a group of order two.
We show that $\mathcal{D}(\mathrm{M}_{p+q})$ is invariant under $\gamma$. In block form, the reflection reads
$\gamma:\begin{pmatrix}A&B\\ D&C\end{pmatrix}\mapsto\begin{pmatrix}A&-B\\ -D&C\end{pmatrix}\,,$
where $A\in\mathrm{M}_{p}$, $B\in\mathbb{C}^{p\times q}$, $C\in\mathrm{M}_{q}$, and $D\in\mathbb{C}^{q\times p}$. The space of hermitian matrices is invariant under $\gamma$, which restricts to
$\mathcal{H}(\mathrm{M}_{p+q})\to\mathcal{H}(\mathrm{M}_{p+q})\,,\quad M=\begin{pmatrix}A&B\\ B^{\ast}&C\end{pmatrix}\mapsto M^{\prime}=\begin{pmatrix}A&-B\\ -B^{\ast}&C\end{pmatrix}\,,$
where $A\in\mathcal{H}(\mathrm{M}_{p})$, $B\in\mathbb{C}^{p\times q}$, and $C\in\mathcal{H}(\mathrm{M}_{q})$. By Rem. 6.2, the map $M\mapsto M^{\prime}$ preserves the positive semidefiniteness since $M$ and $M^{\prime}$ have the same diagonal blocks and both off-diagonal blocks differ in a sign, so that $M^{\prime}/A=M/A$ holds. The map $\gamma$ preserves the trace. The set of fixed points $\mathrm{M}_{p}\oplus\mathrm{M}_{q}=\{A\in\mathrm{M}_{p+q}:\gamma(A)=A\}$ is a *-subalgebra of $\mathrm{M}_{p+q}$. ∎
Corollary 6.4.
The orthogonal projection $\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathrm{M}_{n_{1}}\oplus\dots\oplus\mathrm{M}_{n_{k}})$ is open, where $n=n_{1}+\dots+n_{k}$.
Proof.
Proceeding by induction, we observe that the orthogonal projection $\mathcal{D}(\mathrm{M}_{n_{1}})\to\mathcal{D}(\mathrm{M}_{n_{1}})$ is of course open. Let $k\geq 1$ and assume that the orthogonal projection
$\pi_{k}:\mathcal{D}(\mathrm{M}_{n_{1}+\dots+n_{k}})\to\mathcal{D}(\mathrm{M}_{n_{1}}\oplus\dots\oplus\mathrm{M}_{n_{k}})$
is open. The orthogonal projection $\mathrm{M}_{n_{1}+\dots+n_{k+1}}\to\mathrm{M}_{n_{1}}\oplus\dots\oplus\mathrm{M}_{n_{k+1}}$ factors into the orthogonal projections
$\mathrm{M}_{n_{1}+\dots+n_{k+1}}$ $\to\mathrm{M}_{n_{1}+\dots+n_{k}}\oplus\mathrm{M}_{n_{k+1}}$
$\text{and}\quad\mathrm{M}_{n_{1}+\dots+n_{k}}\oplus\mathrm{M}_{n_{k+1}}$ $\to\mathrm{M}_{n_{1}}\oplus\dots\oplus\mathrm{M}_{n_{k}}\oplus\mathrm{M}_{n_{k+1}}\,.$
Hence, by Lemma 2.5, the map $\pi_{k+1}$ factors into the orthogonal projections
$\mathcal{D}(\mathrm{M}_{n_{1}+\dots+n_{k+1}})$ $\to\mathcal{D}(\mathrm{M}_{n_{1}+\dots+n_{k}}\oplus\mathrm{M}_{n_{k+1}})$ (6.3)
$\text{and}\quad\mathcal{D}(\mathrm{M}_{n_{1}+\dots+n_{k}}\oplus\mathrm{M}_{n_{k+1}})$ $\to\mathcal{D}(\mathrm{M}_{n_{1}}\oplus\dots\oplus\mathrm{M}_{n_{k}}\oplus\mathrm{M}_{n_{k+1}})\,.$ (6.4)
The map (6.3) is open by Prop. 6.3. Lemma 3.7 shows that (6.4) equals
$\mathcal{D}(\mathrm{M}_{n_{1}+\dots+n_{k}})\oplus_{\mathrm{c}}\mathcal{D}(\mathrm{M}_{n_{k+1}})$ $\to\mathcal{D}(\mathrm{M}_{n_{1}}\oplus\dots\oplus\mathrm{M}_{n_{k}})\oplus_{\mathrm{c}}\mathcal{D}(\mathrm{M}_{n_{k+1}})\,,$
which is open by Prop. 3.2 (for $m=2$) and by the induction hypothesis. Being the composition of two open maps, $\pi_{k+1}$ is open. ∎
Proposition 6.5.
The orthogonal projection $\mathcal{D}(\bigoplus_{i=1}^{k}\mathrm{M}_{q})\to\mathcal{D}(\mathrm{M}_{q}\otimes\mathds{1}_{k})$ is open.
Proof.
Lemma 6.1 proves the claim when $\mathcal{R}:=\bigoplus_{i=1}^{k}\mathrm{M}_{q}$ and $\mathcal{A}:=\mathrm{M}_{q}\otimes\mathds{1}_{k}$. The convex set $\mathcal{D}(\mathcal{R})$ is stable by Rem. 5.1 c) and equals the $k$-fold direct convex sum $\mathcal{D}(\mathcal{R})=\mathcal{D}(\mathrm{M}_{q})\oplus_{\mathrm{c}}\cdots\oplus_{\mathrm{c}}\mathcal{D}(\mathrm{M}_{q})$ of $\mathcal{D}(\mathrm{M}_{q})$ by Lemma 3.4. Hence, the cyclic permutation $\gamma=(1,\dots,k)$ defines the orthogonal transformation
$\gamma:\mathcal{D}(\mathcal{R})\to\mathcal{D}(\mathcal{R})\,,\quad(\sigma_{1},\dots,\sigma_{k})\mapsto(\sigma_{\gamma^{-1}(1)},\dots,\sigma_{\gamma^{-1}(k)})\,,$
which generates a group of order $k$. Clearly, $\mathcal{D}(\mathcal{R})$ is invariant under $\gamma$ and $\mathcal{A}=\{A\in\mathcal{R}:\gamma(A)=A\}$ is a *-subalgebra of $\mathrm{M}_{kq}$. ∎
Proof of Thm. 1.2.
Since $\mathcal{A}$ is a *-subalgebra of $\mathrm{M}_{n}$, there exists a unitary $n\times n$ matrix $U$ such that $U\mathcal{A}U^{\ast}=\bigoplus_{i=1}^{m}\mathcal{A}_{i}$, where $\mathcal{A}_{i}:=\mathrm{M}_{q_{i}}\otimes\mathds{1}_{k_{i}}$ for every $i=1,\dots,m$, and $q_{1}k_{1}+\dots+q_{m}k_{m}=n$, see [21, Thm. 5.6].
As $\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathrm{M}_{n})$, $\rho\mapsto U\rho U^{\ast}$ is a homeomorphism, it suffices to prove that the orthogonal projection $\pi:\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\bigoplus_{i=1}^{m}\mathcal{A}_{i})$ is open. The orthogonal projection $\mathrm{M}_{n}\to\bigoplus_{i=1}^{m}\mathcal{A}_{i}$ factors into the orthogonal projections
$\mathrm{M}_{n}$ $\textstyle\to\mathcal{B}_{1}\oplus\dots\oplus\mathcal{B}_{m}$
$\text{and}\quad\mathcal{B}_{1}\oplus\dots\oplus\mathcal{B}_{m}$ $\textstyle\to\mathcal{A}_{1}\oplus\dots\oplus\mathcal{A}_{m}\,,$
where $\mathcal{B}_{i}:=\bigoplus_{j=1}^{k_{i}}\mathrm{M}_{q_{i}}$, $i=1,\dots,m$. By Lemma 2.5, the map $\pi$ factors into
$\mathcal{D}(\mathrm{M}_{n})$ $\to\mathcal{D}(\mathcal{B}_{1}\oplus\dots\oplus\mathcal{B}_{m})\,,$ (6.5)
$\text{and}\quad\mathcal{D}(\mathcal{B}_{1}\oplus\dots\oplus\mathcal{B}_{m})$ $\to\mathcal{D}(\mathcal{A}_{1}\oplus\dots\oplus\mathcal{A}_{m})\,.$ (6.6)
The map (6.5) is open by Coro. 6.4 (for $k=k_{1}+\dots+k_{m}$). Lemma 3.7 shows that the map (6.6) equals
$\mathcal{D}(\mathcal{B}_{1})\oplus_{\mathrm{c}}\cdots\oplus_{\mathrm{c}}\mathcal{D}(\mathcal{B}_{m})$ $\to\mathcal{D}(\mathcal{A}_{1})\oplus_{\mathrm{c}}\cdots\oplus_{\mathrm{c}}\mathcal{D}(\mathcal{A}_{m})\,,$
which is open by Prop. 3.2 and Prop. 6.5. In conclusion, $\pi$ is open as it is a composition of two open maps. ∎
7. Real *-subalgebras of $\mathrm{M}_{n}$
Every real *-subalgebra of $\mathrm{M}_{n}$ is *-isomorphic [21, Thm. 5.22] to a direct sum of algebras of real, complex, and quaternionic $q$-by-$q$-matrices of various sizes $q$. We are here interested in the algebra $\mathrm{M}_{n}(\mathbb{R})$ of real $n\times n$ matrices.
Proposition 7.1.
The orthogonal projection $\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathrm{M}_{n}(\mathbb{R}))$ is open.
Proof.
Lemma 6.1 proves the claim when $\mathcal{R}:=\mathrm{M}_{n}$ and $\mathcal{A}:=\mathrm{M}_{n}(\mathbb{R})$. The set of density matrices $\mathcal{D}(\mathrm{M}_{n})$ is stable by Rem. 5.1 b). The reflection $\gamma:\mathrm{M}_{n}\to\mathrm{M}_{n}$ at the real subspace $\mathrm{M}_{n}(\mathbb{R})$ generates a group of order two.
The algebra $\mathrm{M}_{n}$ is the orthogonal direct sum $\mathrm{M}_{n}=\mathrm{M}_{n}(\mathbb{R})\oplus\mathrm{i}\,\mathrm{M}_{n}(\mathbb{R})$ and the reflection at the real subspace $\mathrm{M}_{n}(\mathbb{R})$ reads
$\gamma:\mathrm{M}_{n}\to\mathrm{M}_{n}\,,\quad A+\mathrm{i}B\mapsto A-\mathrm{i}B\,,\quad A,B\in\mathrm{M}_{n}(\mathbb{R})\,.$
The orthogonal transformation $\gamma$ preserves the space of hermitian matrices, which is the orthogonal direct sum
$\mathcal{H}(\mathrm{M}_{n})=\mathrm{Sym}_{n}(\mathbb{R})\oplus\mathrm{i}\,\mathrm{Skew}_{n}(\mathbb{R})$
of the space of real symmetric matrices
$\mathrm{Sym}_{n}(\mathbb{R})$ :=$:=\{A\in\mathrm{M}_{n}(\mathbb{R})\colon A{}^{\mathrm{T}}=A\}=\mathcal{H}(\mathrm{M}_{n}(\mathbb{R}))$
and the space of skew-symmetric matrices
$\mathrm{Skew}_{n}(\mathbb{R})$ :=$:=\{A\in\mathrm{M}_{n}(\mathbb{R})\colon A{}^{\mathrm{T}}=-A\}\,.$
We prove that $\gamma$ preserves the trace and the positive semidefiniteness on the space of hermitian matrices. Let $A,C\in\mathrm{Sym}_{n}(\mathbb{R})$ and $B,D\in\mathrm{Skew}_{n}(\mathbb{R})$. Then $\operatorname{tr}(A+\mathrm{i}B)=\operatorname{tr}(A)$ shows that the trace is preserved. It is well known that a matrix is positive semidefinite if and only if its inner product with the square of every hermitian matrix is nonnegative. Thus
$\langle A-\mathrm{i}B,(C+\mathrm{i}D)^{2}\rangle$ $=\langle A+\mathrm{i}B,(C-\mathrm{i}D)^{2}\rangle$
shows that $A-\mathrm{i}B$ is positive semidefinite if $A+\mathrm{i}B$ is positive semidefinite. Clearly, $\mathrm{M}_{n}(\mathbb{R})=\{A\in\mathrm{M}_{n}:\gamma(A)=A\}$ is a real *-subalgebra of $\mathrm{M}_{n}$. ∎
As the orthogonal projection $\mathrm{M}_{n}\to\mathrm{M}_{n}(\mathbb{R})$ is the entrywise real part, we denote by $\mathrm{Re}:\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathrm{M}_{n}(\mathbb{R}))$ the orthogonal projection of $\mathcal{D}(\mathrm{M}_{n})$ onto $\mathcal{D}(\mathrm{M}_{n}(\mathbb{R}))$.
Example 7.2.
We consider the chain $\mathrm{M}_{3}\supset\mathrm{M}_{2}\oplus\mathrm{M}_{1}\supset\mathrm{M}_{2}(\mathbb{R})\oplus\mathrm{M}_{1}(\mathbb{R})$ of real *-subalgebras of $\mathrm{M}_{3}$. The orthogonal projections
$\mathrm{M}_{3}\longrightarrow\mathrm{M}_{2}\oplus\mathrm{M}_{1}\longrightarrow\mathrm{M}_{2}(\mathbb{R})\oplus\mathrm{M}_{1}(\mathbb{R})$
restrict by Lemma 2.5 to
$\mathcal{D}(\mathrm{M}_{3})\stackrel{{\scriptstyle\pi_{1}}}{{\longrightarrow}}\mathcal{D}(\mathrm{M}_{2}\oplus\mathrm{M}_{1})\stackrel{{\scriptstyle\pi_{2}}}{{\longrightarrow}}\mathcal{D}(\mathrm{M}_{2}(\mathbb{R})\oplus\mathrm{M}_{1}(\mathbb{R}))\,.$
The map $\pi_{1}$ is open by Prop. 6.3. By Lemma 3.7, the map $\pi_{2}$ is the direct convex sum
$\mathrm{Re}\oplus_{\mathrm{c}}\operatorname{id}:\mathcal{D}(\mathrm{M}_{2})\oplus_{\mathrm{c}}\{1\}\to\mathcal{D}(\mathrm{M}_{2}(\mathbb{R}))\oplus_{\mathrm{c}}\{1\}$
of $\mathrm{Re}:\mathcal{D}(\mathrm{M}_{2})\to\mathcal{D}(\mathrm{M}_{2}(\mathbb{R}))$ and the identity map $\operatorname{id}:\{1\}\to\{1\}$. The map $\mathrm{Re}$ is open by Prop. 7.1, hence $\pi_{2}$ is open by Prop. 3.2. The orthogonal projection $\mathcal{D}(\mathrm{M}_{3})\to\mathcal{D}(\mathrm{M}_{2}(\mathbb{R})\oplus\mathrm{M}_{1}(\mathbb{R}))$ is open, as it is a composition of two open maps.
8. Topology simplified by algebra
Thm. 1.2 can simplify topology problems. Given topological spaces $K,L$, a map $f:K\to L$ is continuous [27] at $x\in K$ if the preimage of every neighborhood of $f(x)$ in $L$ is a neighborhood of $x$ in $K$.
Lemma 8.1.
Let $\mathcal{R}_{1},\mathcal{R}_{2}$ be real matrix systems on $\mathbb{C}^{n}$ such that $\mathcal{R}_{2}\subset\mathcal{R}_{1}$. Let $\pi_{1}:\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathcal{R}_{1})$ and $\pi_{2}:\mathcal{D}(\mathcal{R}_{1})\to\mathcal{D}(\mathcal{R}_{2})$ denote the orthogonal projections and assume the orthogonal projection $\pi_{2}\circ\pi_{1}:\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathcal{R}_{2})$ is open. Let $f:\mathcal{D}(\mathcal{R}_{2})\to T$ be a map to a topological space $T$. Let $\rho\in\mathcal{D}(\mathcal{R}_{1})$.
  1. a)
    The map $f\circ\pi_{2}:\mathcal{D}(\mathcal{R}_{1})\to T$ is open at $\rho$ if and only if $f:\mathcal{D}(\mathcal{R}_{2})\to T$ is open at $\pi_{2}(\rho)$.
  2. b)
    The map $f\circ\pi_{2}:\mathcal{D}(\mathcal{R}_{1})\to T$ is continuous at $\rho$ if and only if the map $f:\mathcal{D}(\mathcal{R}_{2})\to T$ is continuous at $\pi_{2}(\rho)$.
Proof.
The orthogonal projection $\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathcal{R}_{2})$ equals indeed $\pi_{2}\circ\pi_{1}$ by Lemma 2.5.
We begin with the implication $\Rightarrow$” of a). If $\mathcal{N}_{2}\subset\mathcal{D}(\mathcal{R}_{2})$ is a neighborhood of $\pi_{2}(\rho)$, then
$f(\mathcal{N}_{2})=(f\circ\pi_{2})\circ\pi_{2}^{-1}(\mathcal{N}_{2})$
is a neighborhood of $f(\pi_{2}(\rho))$ because $\pi_{2}$ is continuous and $f\circ\pi_{2}$ is open at $\rho$. Regarding the implication $\Leftarrow$”, we choose a neighborhood $\mathcal{N}_{1}\subset\mathcal{D}(\mathcal{R}_{1})$ of $\rho$. Then
$f\circ\pi_{2}(\mathcal{N}_{1})=f\circ(\pi_{2}\circ\pi_{1})\circ\pi_{1}^{-1}(\mathcal{N}_{1})$
is a neighborhood of $f(\pi_{2}(\rho))$ because $\pi_{1}$ is continuous, $\pi_{2}\circ\pi_{1}$ is open, and $f$ is open at $\pi_{2}(\rho)$.
To prove b) we choose a neighborhood $\mathcal{N}_{T}\subset T$ of $f(\pi_{2}(\rho))$. Regarding the implication $\Rightarrow$”, the preimage
$f^{-1}(\mathcal{N}_{T})=(\pi_{2}\circ\pi_{1})\circ\pi_{1}^{-1}\circ(f\circ\pi_{2})^{-1}(\mathcal{N}_{T})$
is a neighborhood of $\pi_{2}(\rho)$, because $f\circ\pi_{2}$ is continuous at $\rho$, the map $\pi_{1}$ is continuous, and $\pi_{2}\circ\pi_{1}$ is open. Regarding the implication $\Leftarrow$”, the preimage
$(f\circ\pi_{2})^{-1}(\mathcal{N}_{T})=\pi_{2}^{-1}\circ f^{-1}(\mathcal{N}_{T})$
is a neighborhood of $\rho$, as $f$ is continuous at $\pi_{2}(\rho)$, and $\pi_{2}$ is continuous. ∎
Remark 8.2 (Simplifying openness problems).
Let $\pi:\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathcal{R})$ be the orthogonal projection to a real matrix system $\mathcal{R}$ on $\mathbb{C}^{n}$ and let $\mathcal{A}$ be a *-subalgebra of $\mathrm{M}_{n}$ such that $\mathcal{R}\subset\mathcal{A}$. Then $\pi=\pi_{\mathcal{R}}\circ\pi_{\mathcal{A}}$ factors into the orthogonal projections $\pi_{\mathcal{A}}:\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathcal{A})$ and $\pi_{\mathcal{R}}:\mathcal{D}(\mathcal{A})\to\mathcal{D}(\mathcal{R})$. For every $\rho\in\mathcal{D}(\mathrm{M}_{n})$ the map $\pi$ is open at $\rho$ if and only if $\pi_{\mathcal{R}}$ is open at $\pi_{\mathcal{A}}(\rho)$.
As a word of explication, the map $\pi$ factors by Lemma 2.5. The claim follows from Lemma 8.1 a), by letting $\mathcal{R}_{1}:=\mathrm{M}_{n}$ and $\mathcal{R}_{2}:=\mathcal{A}$, and by taking $\pi_{\mathcal{R}}$ as the map $f:\mathcal{D}(\mathcal{R}_{2})\to T$. The assumptions of the lemma are met since $\pi_{\mathcal{A}}$ is open by Thm. 1.2.
The following two examples are concerned with openness. Continuity is a topic of Sec. 9 below.
Example 8.3.
Let $\mathcal{R}:=\operatorname{span}_{\mathbb{R}}(\mathds{1}_{3},X\oplus 1,Z\oplus 0)$.
  1. a)
    Rem. 8.2 solves the openness problem of the orthogonal projection $\pi:\mathcal{D}(\mathrm{M}_{3})\to\mathcal{D}(\mathcal{R})$ to $\mathcal{R}$, based on the solution of the openness problem of $\pi_{\mathcal{R}}:\mathcal{D}(\mathcal{A})\to\mathcal{D}(\mathcal{R})$, where $\mathcal{A}:=\mathrm{M}_{2}\oplus\mathrm{M}_{1}$. By Ex. 3.5 a), the map $\pi_{\mathcal{R}}$ is not open at any point in the half-open segment $\mathcal{G}_{0}=\big{(}\ket{+}\!\!\bra{+}\oplus 0,0\oplus 1\big{]}\subset\mathcal{D}(\mathcal{A})$ and open at every point in the complement. By Rem. 8.2, the map $\pi:\mathcal{D}(\mathrm{M}_{3})\to\mathcal{D}(\mathcal{R})$ is not open at any point in $\pi_{\mathcal{A}}^{-1}(\mathcal{G}_{0})$ and open at every point in the complement.
  2. b)
    To describe $\pi_{\mathcal{A}}^{-1}(\mathcal{G}_{0})$, we study the fibers of the orthogonal projection
    $\mathcal{H}(\mathrm{M}_{3})\to\mathcal{H}(\mathcal{A})\,,\quad\begin{pmatrix}A&\ket{\varphi}\\ \bra{\varphi}&c\end{pmatrix}\mapsto\begin{pmatrix}A&0\\ 0&c\end{pmatrix}\,,$
    where $A\in\mathcal{H}(\mathrm{M}_{2})$, |$\ket{\varphi}\in\mathbb{C}^{2}\cong\mathbb{C}^{2\times 1}$, and $c\in\mathbb{R}\cong\mathcal{H}(\mathrm{M}_{1})$. Every point in $\mathcal{G}=\big{[}\ket{+}\!\!\bra{+}\oplus 0,0\oplus 1\big{]}$ is of the form $(1-\lambda)\ket{+}\!\!\bra{+}\oplus\lambda$ for some $\lambda\in[0,1]$. Using the generalized Schur complement (Rem. 6.2), one verifies that the fiber of $\pi_{\mathcal{A}}$ over this point is the set of all matrices
    $\rho(\lambda,z):=\begin{pmatrix}(1-\lambda)\ket{+}\!\!\bra{+}&z\ket{+}\\ \overline{z}\bra{+}&\lambda\end{pmatrix}\,,\quad z\in\mathbb{C}\,,\quad|z|^{2}\leq\lambda(1-\lambda)\,,$ (8.1)
    where $|z|$ denotes the absolute value of $z\in\mathbb{C}$. In conclusion, the orthogonal projection $\pi:\mathcal{D}(\mathrm{M}_{3})\to\mathcal{D}(\mathcal{R})$ is not open at any point of
    $\pi_{\mathcal{A}}^{-1}(\mathcal{G}_{0})=\{\rho(\lambda,z)\colon z\in\mathbb{C},|z|^{2}\leq\lambda(1-\lambda),\lambda\in(0,1]\}$
    and open at every point in the complement.
  3. c)
    We verify a claim made in Ex. 4.6. The segment $\mathcal{G}=\big{[}\ket{+}\!\!\bra{+}\oplus 0,0\oplus 1\big{]}$ is the fiber of $\pi_{\mathcal{R}}$ over $M=\frac{1}{2}(\ket{+}\!\!\bra{+}\oplus 1)\in\mathcal{D}(\mathcal{R})$ by Ex. 3.5 b), so $\pi_{\mathcal{A}}^{-1}(\mathcal{G})$ is the fiber of $\pi=\pi_{\mathcal{R}}\circ\pi_{\mathcal{A}}$ over $M$. As recalled in part a) above, the map $\pi$ is open at |++|$\ket{+}\!\!\bra{+}\oplus 0$ but not open at any point of $\pi_{\mathcal{A}}^{-1}(\mathcal{G}_{0})$. We now observe that $\pi$ is not open at any point in the relative interior of $\pi^{-1}(M)$, as we have the chain of inclusions
[…]
    $\textstyle=\operatorname{ri}(\pi_{\mathcal{R}}^{-1}(M))=\operatorname{ri}(\mathcal{G})=\mathcal{G}_{0}\setminus\{0\oplus 1\}\subset\mathcal{G}_{0}\,,$
    whose first equality holds by [45, Thm. 6.6].
Example 8.4.
The recipe of Rem. 8.2 helps analyze the openness of the orthogonal projection $\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathcal{R}(P,Q))$ to the real matrix system
$\mathcal{R}(P,Q)=\operatorname{span}_{\mathbb{R}}(\mathds{1}_{n},P,Q)$
generated by two orthogonal projections $P,Q\in\mathrm{M}_{n}$, that is to say, matrices satisfying $P=P^{2}=P^{\ast}$ and $Q=Q^{2}=Q^{\ast}$. It is well known that the matrix system $\mathcal{R}(P,Q)$ is included in a surprisingly small *-subalgebra of $\mathrm{M}_{n}$, see Coro. 2.2 in the survey [8] by Böttcher and Spitkovsky, and the references therein. More precisely, there exists a unitary $n\times n$ matrix $U$, nonnegative integers $m_{1}\leq 4$ and $m_{2}$, and positive integers $k_{i}$, $i=1,\ldots,m_{1}$ satisfying $k_{1}+\dots+k_{m_{1}}+2m_{2}=n$, such that $\mathcal{R}:=U\mathcal{R}(P,Q)U^{\ast}$ is included in
$\mathcal{A}:=\big{(}\bigoplus_{i=1}^{m_{1}}\mathrm{M}_{1}\otimes\mathds{1}_{k_{i}}\big{)}\oplus\big{(}\bigoplus_{j=1}^{m_{2}}\mathrm{M}_{2}\big{)}\,.$
Since $\rho\mapsto U\rho U^{\ast}$ is a homeomorphism of $\mathcal{D}(\mathrm{M}_{n})$, the openness problems of the orthogonal projections $\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathcal{R}(P,Q))$ and $\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathcal{R})$ are equivalent. The second one is substantially simplified by the method of Rem. 8.2 as $\mathcal{D}(\mathcal{A})$ has a rather simple shape. It is the direct convex sum of several three-dimensional Euclidean balls and a simplex of dimension at most three by Lemma 3.4 and Ex. 2.3. This observation should also simplify the strong continuity problem for the numerical range $W(P+\mathrm{i}Q)$.
9. Continuity in quantum information theory
We discuss continuity problems of entropic inference maps and of measures of correlation. We assume that $\mathcal{R}$ is a real matrix system on $\mathbb{C}^{n}$ and that, without loss of generality (see Rem. 2.4), we have $\mathcal{R}\subset\mathcal{H}(\mathrm{M}_{n})$.
Example 9.1 (Maximum entropy inference I).
The purpose of the maximum entropy inference method is to update a prior probability distribution if new information becomes available in the form of constraints that specify a set of possible posterior probability distributions. The preferred posterior is that which minimizes the relative entropy from the prior subject to the available constraints, see Chap. 8 in Caticha’s book [13], and the references therein. Whereas the maximum entropy method has a sound axiomatic basis, the analogous inference method for quantum states lacks an axiomatic foundation and remains a controversial topic [3].
Analytic properties of quantum mechanical inference are interesting despite unsolved foundational issues. Notably, quantum inference maps have discontinuities [55] that aroused interest in theoretical physics [14, 26, 36, 56].
Linear constraints [^4] Linear constraints can be defined in terms of expectation values. Let $A_{1},\dots,A_{k}$ be hermitian $n\times n$ matrices such that $\mathcal{R}=\operatorname{span}_{\mathbb{R}}(\mathds{1}_{n},A_{1},\dots,A_{k})$. The observables represented by $A_{1},\dots,A_{k}$ have the expectation values $v(\rho)=(\langle\rho,A_{1}\rangle,\dots,\langle\rho,A_{k}\rangle){}^{\mathrm{T}}$ if $\rho\in\mathcal{D}(\mathrm{M}_{n})$ is the system state [7]. The fiber $\pi^{-1}(\sigma)$ over $\sigma\in\mathcal{D}(\mathcal{R})$ is the set of $\rho\in\mathcal{D}(\mathrm{M}_{n})$ whose expectation values are $v(\rho)=v(\sigma)$, see Lemma 4.1. on $\mathcal{D}(\mathrm{M}_{n})$ are defined by the orthogonal projection
$\pi:\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathcal{R})\,.$
The relative entropy $S:\mathcal{D}(\mathrm{M}_{n})\times\mathcal{D}(\mathrm{M}_{n})\to[0,+\infty]$ is an asymmetric distance. It is defined as $S(\rho_{1},\rho_{2}):=\operatorname{tr}[\rho_{1}(\log(\rho_{1})-\log(\rho_{2}))]$ if the range of $\rho_{1}$ is included in the range of $\rho_{2}$ and as $S(\rho_{1},\rho_{2}):=+\infty$ otherwise, for all $\rho_{1},\rho_{2}\in\mathcal{D}(\mathrm{M}_{n})$. Let $\tau\in\mathcal{D}(\mathrm{M}_{n})$, the prior, be a density matrix of maximal rank $n$. Then
$\phi_{\tau}:\mathcal{D}(\mathrm{M}_{n})\to\mathbb{R}\,,\quad\rho\mapsto-S(\rho,\tau)$
is continuous and strictly concave. So the maximum entropy inference map
$\Psi_{\tau}:\mathcal{D}(\mathcal{R})\to\mathcal{D}(\mathrm{M}_{n})\,,\quad\sigma\mapsto\operatorname{argmax}_{\rho\in\pi^{-1}(\sigma)}\phi_{\tau}(\rho)$ (9.1)
is well defined, see [52, Def. 1.1] and the references therein.
We quote [52, Thm. 4.9] under the assumptions of Ex. 9.1.
Theorem 9.2.
For any $\sigma\in\mathcal{D}(\mathcal{R})$, the inference map $\Psi_{\tau}:\mathcal{D}(\mathcal{R})\to\mathcal{D}(\mathrm{M}_{n})$ is continuous at $\sigma$ if and only if the orthogonal projection $\pi:\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathcal{R})$ is open at $\Psi_{\tau}(\sigma)$.
Example 9.3 (Maximum entropy inference II).
We resume Ex. 9.1. Thm. 9.2 translates the continuity problem of the inference map $\Psi_{\tau}$ into an openness problem of the orthogonal projection $\pi$. Whether Rem. 8.2 can simplify this openness problem depends on the location of the posteriors.
The range of $\Psi_{\tau}$ can be described explicitly. The manifold
$\mathcal{E}_{\tau}(\mathcal{R}):=\big{\{}\frac{e^{\log(\tau)+A}}{\operatorname{tr}(e^{\log(\tau)+A})}\mid A\in\mathcal{R}\big{\}}$ (9.2)
is known as a Gibbsian family or exponential family, see [55, 38, 52, 43] and the references therein (and recall $\mathcal{R}=\mathcal{H}(\mathcal{R})$). The manifold $\mathcal{E}_{\tau}(\mathcal{R})$ is included in $\Psi_{\tau}(\mathcal{D}(\mathcal{R}))$. But not just that, the set $\Psi_{\tau}(\mathcal{D}(\mathcal{R}))$ of maximum entropy inference states equals
$\{\rho_{1}\in\mathcal{D}(\mathrm{M}_{n}):\inf_{\rho_{2}\in\mathcal{E}_{\tau}(\mathcal{R})}S(\rho_{1},\rho_{2})=0\}\,,$ (9.3)
which is a subset of the Euclidean closure of $\mathcal{E}_{\tau}(\mathcal{R})$, see (D5) in [52]. The set (9.3) is called reverse information closure or the rI-closure [19] of $\mathcal{E}(\mathcal{R})$.
Let $\mathcal{A}$ denote the smallest *-subalgebra $\mathcal{A}$ of $\mathrm{M}_{n}$ that contains $\mathcal{R}$.
  1. a)
    Rem. 8.2 is of no avail if $\tau=\mathds{1}_{n}/n$ is the uniform prior. In this case,
    $\phi_{\tau}(\rho)=-S(\rho,\mathds{1}_{n}/n)=S(\rho)-\log(n)\,,\quad\rho\in\mathcal{D}(\mathrm{M}_{n})$
    is the von Neumann entropy $S(\rho):=-\operatorname{tr}[\rho\log(\rho)]$ up to a constant. By functional calculus, the exponential family $\mathcal{E}_{\tau}(\mathcal{R})$ is included in $\mathcal{A}$ and so is the set of inference states $\Psi_{\tau}(\mathcal{D}(\mathcal{R}))$ as per the formula (9.3), because $\mathcal{A}$ is closed. Rem. 8.2 is of no help since all posteriors lie in $\mathcal{A}$ anyways. The same conclusion is true for every prior $\tau$ in $\mathcal{A}$ replacing $\mathds{1}_{n}/n$.
  2. b)
    If the prior $\tau$ lies outside of $\mathcal{A}$ then the exponential family $\mathcal{E}_{\tau}(\mathcal{R})$ is disjoint from $\mathcal{D}(\mathcal{A})$, again by functional calculus. Rem. 8.2 brings an advantage if the orthogonal projection $\pi_{\mathcal{R}}:\mathcal{D}(\mathcal{A})\to\mathcal{D}(\mathcal{R})$ is easier to analyze than $\pi:\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathcal{R})$. An example is given in Ex. 8.4 where $\mathcal{A}$ has a simple structure and a much smaller dimension than $\mathrm{M}_{n}$.
  3. c)
    The real matrix system $\mathcal{R}:=\operatorname{span}_{\mathbb{R}}(\mathds{1}_{3},X\oplus 1,Z\oplus 0)$ in Ex. 8.3 has the property that the orthogonal projection $\pi:\mathcal{D}(\mathrm{M}_{3})\to\mathcal{D}(\mathcal{R})$ is not open anywhere in the relative interior of the fiber over $M:=\frac{1}{2}(\ket{+}\!\!\bra{+}\oplus 1)$ and open at every point in the complement of that fiber. For every $\sigma\in\mathcal{D}(\mathcal{R})$ and prior $\tau$, the posterior $\Psi_{\tau}(\sigma)$ is contained in the relative interior of the fiber $\pi^{-1}(\sigma)$ over every $\sigma$ by Coro. 5.7 and Lemma 5.8 in [52]. Thm. 9.2 then shows that $\Psi_{\tau}$ is discontinuous at $M$ and continuous at every point different from $M$ (for all priors $\tau$).
Example 9.4 (Maximum entropy inference III).
An important example from physics is the real matrix system of local Hamiltonians [14, 58].
Every unit $i\in\Omega:=\{1,2,\ldots,N\}$ of an $N$-qubit system is associated with a copy $\mathcal{A}_{i}$ of the algebra $\mathrm{M}_{2}$. The subsystem with units in a subset $\nu\subset\Omega$ is associated with the tensor product algebra $\mathcal{A}_{\nu}:=\bigotimes_{i\in\nu}\mathcal{A}_{i}$, whose identity we denote by $\mathds{1}_{\nu}$. We have $\mathcal{A}_{\Omega}=\mathrm{M}_{n}$ for $n=2^{N}$. The algebra $\mathcal{A}_{\nu}$ embeds into $\mathcal{A}_{\Omega}$ via the map $\mathcal{A}_{\nu}\to\mathcal{A}_{\Omega}$, $A\mapsto A\otimes\mathds{1}_{\bar{\nu}}$, where $\bar{\nu}=\Omega\setminus\nu$ is the complement of $\nu$. Let $\mathfrak{g}$ be a family of subsets of $\Omega$. A $\mathfrak{g}$-local Hamiltonian is a hermitian matrix in $\mathcal{A}_{\Omega}$ of the form
$\sum_{\nu\in\mathfrak{g}}A_{\nu}\otimes\mathds{1}_{\bar{\nu}}\,,\quad A_{\nu}\in\mathcal{H}(\mathcal{A}_{\nu})\,,\quad\nu\in\mathfrak{g}\,.$ (9.4)
We denote the real matrix system of all $\mathfrak{g}$-local Hamiltonians by $\mathcal{R}_{\mathfrak{g}}$ and the orthogonal projection by $\pi_{\mathfrak{g}}:\mathcal{D}(\mathcal{A}_{\Omega})\to\mathcal{D}(\mathcal{R}_{\mathfrak{g}})$.
The partial trace $\operatorname{Tr}_{\bar{\nu}}:\mathcal{A}_{\Omega}\to\mathcal{A}_{\nu}$ is the adjoint of the embedding $\mathcal{A}_{\nu}\to\mathcal{A}_{\Omega}$ and satisfies $\langle A\otimes\mathds{1}_{\bar{\nu}},B\rangle=\langle A,\operatorname{Tr}_{\bar{\nu}}(B)\rangle$ for every $A\in\mathcal{A}_{\nu}$, $B\in\mathcal{A}_{\Omega}$. The partial trace $\operatorname{Tr}_{\bar{\nu}}(\rho)$ of $\rho\in\mathcal{D}(\mathcal{A}_{\Omega})$ is a density matrix of $\mathcal{A}_{\nu}$ called reduced density matrix. Let
$\operatorname{red}_{\mathfrak{g}}:\mathcal{A}_{\Omega}\to\prod_{\nu\in\mathfrak{g}}\mathcal{A}_{\nu}\,,\quad A\mapsto[\operatorname{Tr}_{\bar{\nu}}(A)]_{\nu\in\mathfrak{g}}$ (9.5)
denote the map from $\mathcal{A}_{\Omega}$ to the cartesian product of the algebras $(\mathcal{A}_{\nu})_{\nu\in\mathfrak{g}}$ that assigns partial traces to the subsystems specified by the members of $\mathfrak{g}$.
Linear constraints on $\mathcal{D}(\mathcal{A}_{\Omega})$ have been defined in terms of reduced density matrices, see [38, 14] and [58, Sec. 1.4.2]. This is formalized in the following diagram, which commutes by formula (19) in [54]. (Obvious restrictions of $\operatorname{red}_{\mathfrak{g}}$ are omitted.)
${\mathcal{D}(\mathcal{A}_{\Omega})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{\mathfrak{g}}}$$\scriptstyle{\operatorname{red}_{\mathfrak{g}}}$${\operatorname{red}_{\mathfrak{g}}[\mathcal{D}(\mathcal{A}_{\Omega})]}$${\mathcal{D}(\mathcal{R}_{\mathfrak{g}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\operatorname{red}_{\mathfrak{g}}}$
The commutative diagram shows that the fiber $\pi_{\mathfrak{g}}^{-1}(\sigma)$ over $\sigma\in\mathcal{D}(\mathcal{R}_{\mathfrak{g}})$ is the set of all $\rho\in\mathcal{D}(\mathcal{A}_{\Omega})$ such that $\operatorname{red}_{\mathfrak{g}}(\rho)=\operatorname{red}_{\mathfrak{g}}(\sigma)$. Thm. 9.2 proves that the pullback $\Xi_{\tau}:=\Psi_{\tau}\circ\operatorname{red}_{\mathfrak{g}}^{-1}$ of the inference map $\Psi_{\tau}$ under $\operatorname{red}_{\mathfrak{g}}^{-1}$ is continuous at $(\rho_{\nu})_{\nu\in\mathfrak{g}}\in\operatorname{red}_{\mathfrak{g}}[\mathcal{D}(\mathcal{A}_{\Omega})]$ if and only if $\operatorname{red}_{\mathfrak{g}}$ is open at $\Xi_{\tau}[(\rho_{\nu})_{\nu\in\mathfrak{g}}]\in\mathcal{D}(\mathcal{A}_{\Omega})$.
A discontinuity of $\Xi_{\tau}$ (with uniform prior $\tau$) was reported for $N=3$ qubits and $\mathfrak{g}:=\{\{1,2\},\{2,3\},\{3,1\}\}$ at $(\rho_{\nu})_{\nu\in\mathfrak{g}}\in\operatorname{red}_{\mathfrak{g}}[\mathcal{D}(\mathcal{A}_{\Omega})]$, where
$\rho_{\nu}:=\frac{1}{2}(\ket{00}\!\!\bra{00}+\ket{11}\!\!\bra{11})\,,\quad\nu\in\mathfrak{g}\,.$
This result by Chen et al. [14, Ex. 4] has an interesting interpretation. The map $\operatorname{red}_{\mathfrak{g}}$ being open [^5] The openness of $\operatorname{red}_{\mathfrak{g}}$ at $\rho\in\mathcal{D}(\mathcal{A}_{\Omega})$ is a priori weaker than the continuity of $\Xi_{\tau}$ at $\operatorname{red}_{\mathfrak{g}}(\rho)$. The continuity means that any sufficiently small change of $\operatorname{red}_{\mathfrak{g}}(\rho)$ can be matched by an arbitrarily small change of $\rho$ inside the range of $\Xi_{\tau}$ and not just anywhere within $\mathcal{D}(\mathcal{A}_{\Omega})$. Somewhat surprisingly, the two propositions are equivalent by Thm. 9.2. at $\rho\in\mathcal{D}(\mathcal{A}_{\Omega})$ means that any sufficiently small change of $\operatorname{red}_{\mathfrak{g}}(\rho)$ within $\operatorname{red}_{\mathfrak{g}}[\mathcal{D}(\mathcal{A}_{\Omega})]$ is matched by an arbitrarily small change of $\rho$ within $\mathcal{D}(\mathcal{A}_{\Omega})$. Conversely, if the openness fails, then there are arbitrarily small changes of $\operatorname{red}_{\mathfrak{g}}(\rho)$ that are only matched by changes of $\rho$ beyond some positive threshold (in the metric sense). Loosely speaking, a small change of a subsystem abruptly changes the entire system. Such behavior is associated with phase transitions [14]. It motivates every attempt to study the openness problem of $\pi_{\mathfrak{g}}$. This should be done by tuning the interaction pattern $\mathfrak{g}$ to a concrete system. Whether enclosing $\mathcal{R}_{\mathfrak{g}}$ into a *-subalgebra of $\mathcal{A}_{\Omega}$ could simplify the problem, as suggested in Rem. 8.2, is not yet clarified.
Focusing on the continuity of correlations measures, we begin with theory.
Example 9.5 (Entropy distance I).
Let $\tau:=\mathds{1}_{n}/n$ and let $\mathcal{E}(\mathcal{R}):=\mathcal{E}_{\tau}(\mathcal{R})$ be an exponential family, see Ex. 9.3. The entropy distance from $\mathcal{E}(\mathcal{R})$ is
$d:\mathcal{D}(\mathrm{M}_{n})\to\mathbb{R}\,,\quad\rho_{1}\mapsto\inf_{\rho_{2}\in\mathcal{E}(\mathcal{R})}S(\rho_{1},\rho_{2})\,.$ (9.6)
Using the von Neumann entropy $S=\phi_{\tau}+\log(n)$, we define
$\Psi:\mathcal{D}(\mathcal{R})\to\mathcal{D}(\mathrm{M}_{n})\,,\quad\sigma\mapsto\operatorname{argmax}_{\rho\in\pi^{-1}(\sigma)}S(\rho)\,.$
Clearly, the map $\Psi$ equals the inference map $\Psi_{\tau}$ defined earlier in Ex. 9.1. The entropy distance can be written as the difference
$d(\rho)=S(\Psi\circ\pi(\rho))-S(\rho)\,,\quad\rho\in\mathcal{D}(\mathrm{M}_{n})$ (9.7)
between the value of the von Neumann entropy at $\rho$ and its maximal value on the fiber of $\pi:\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathcal{R})$ that contains $\rho$, see p. 1288 in [52].
Formula (9.7) suggests studying the rI-projection [19]
$\Pi:\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathrm{M}_{n})\,,\quad\Pi=\Psi\circ\pi\,.$
By (9.3), the range of $\Pi$ is the rI-closure of $\mathcal{E}(\mathcal{R})$. We also consider the restriction $d|_{\mathcal{D}(\mathcal{A})}$ of $d$ and the following restriction of $\Pi$.
$\Pi|_{\mathcal{D}(\mathcal{A})}:\mathcal{D}(\mathcal{A})\to\mathcal{D}(\mathcal{A})\,,\quad\rho\mapsto\Pi(\rho)$
Here, $\mathcal{A}$ is real *-subalgebra of $\mathrm{M}_{n}$ that includes $\mathcal{R}$. The map $\Pi|_{\mathcal{D}(\mathcal{A})}$ is well defined because $\Psi(\mathcal{D}(\mathcal{R}))$ is included in $\mathcal{D}(\mathcal{A})$, as observed in Ex. 9.3 a).
Example 9.6 (Entropy distance II).
The exponential family $\mathcal{E}(\mathcal{R}_{\mathfrak{g}})$ of local Hamiltonians (Ex. 9.4) is interesting because the entropy distance from it quantifies many-body correlations. Amari [4] and Ay [6] studied this type of correlation measure in probability theory. Linden et al. [33] introduced it as a difference of von Neumann entropies like formula (9.7), see also [58, Sec. 1.4.2]. And Zhou [59] proved the equivalence of the two representations (9.6) and (9.7) for density matrices of maximal rank $n$, see also [38]; the equivalence is true without rank restrictions as presented in Ex. 9.5.
We quote from [52, Lemma 5.15 and 4.5] under the assumptions of Ex. 9.5.
Lemma 9.7.
Le $\mathcal{A}$ be real *-subalgebra of $\mathrm{M}_{n}$ that includes $\mathcal{R}$.
  1. a)
    For every $\rho\in\mathcal{D}(\mathcal{A})$, the map $\Pi|_{\mathcal{D}(\mathcal{A})}$ is continuous at $\rho$ if and only if $d|_{\mathcal{D}(\mathcal{A})}$ is continuous at $\rho$.
  2. b)
    For every $\sigma\in\mathcal{D}(\mathcal{R})$, the inference map $\Psi$ is continuous at $\sigma$ if and only if $d$ is continuous at every point in the fiber $\pi^{-1}(\sigma)$.
With reference to Ex. 9.5, we simplify the continuity problem of the entropy distance (and of other invariant maps) using *-subalgebras.
Remark 9.8 (Continuity of the rI-projection).
Let $\mathcal{A}$ be a *-subalgebra of $\mathrm{M}_{n}$ such that $\mathcal{R}\subset\mathcal{A}$. As in Rem. 8.2, the map $\pi$ factors into $\pi=\pi_{\mathcal{R}}\circ\pi_{\mathcal{A}}$. Then, for every $\rho\in\mathcal{D}(\mathrm{M}_{n})$ the rI-projection $\Pi$ is continuous at $\rho$ if and only if $\Pi|_{\mathcal{D}(\mathcal{A})}$ is continuous at $\pi_{\mathcal{A}}(\rho)$.
This assertions needs a justification. By its definition, the map $\Pi=\Psi\circ\pi$ is invariant under $\pi_{\mathcal{A}}$ in the sense that $\Pi=\Pi\circ\pi_{\mathcal{A}}$ holds. This shows $\Pi=\Pi|_{\mathcal{D}(\mathcal{A})}\circ\pi_{\mathcal{A}}$. The claim then follows from Lemma 8.1 b), by letting $\mathcal{R}_{1}:=\mathrm{M}_{n}$ and $\mathcal{R}_{2}:=\mathcal{A}$, and by taking $\Pi|_{\mathcal{D}(\mathcal{A})}$ as the map $f:\mathcal{D}(\mathcal{R}_{2})\to T$. The assumptions of the lemma are met since $\pi_{\mathcal{A}}$ is open by Thm. 1.2.
Example 9.9 (Entropy distance III).
Rem. 9.8 helps solve the continuity problem of the entropy distance from $\mathcal{E}(\mathcal{R})$ for $\mathcal{R}:=\operatorname{span}_{\mathbb{R}}(\mathds{1}_{3},X\oplus 1,Z\oplus 0)$, based on a solution for $d|_{\mathcal{A}(\mathcal{D})}$ in the algebra $\mathcal{A}:=\mathrm{M}_{2}(\mathbb{R})\oplus\mathrm{M}_{1}(\mathbb{R})$ that was obtained using asymptotic curvature estimates [52].
Real *-subalgebras are excluded from Rem. 9.8 but the conclusion is still true, as Ex. 7.2 can replace Thm. 1.2 in this remark.
We recall from Ex. 3.6 that the generatrix $\mathcal{G}=[\ket{+}\!\!\bra{+}\oplus 0,0\oplus 1]$ of the cone $\mathcal{D}(\mathcal{A})$ is the fiber of the orthogonal projection $\pi_{\mathcal{R}}:\mathcal{D}(\mathcal{A})\to\mathcal{D}(\mathcal{R})$ over $M:=\frac{1}{2}(\ket{+}\!\!\bra{+}\oplus 1)$. Thm. 5.18 in [52] shows that $d|_{\mathcal{D}(\mathcal{A})}$ is discontinuous at every point in the half-open segment
$\mathcal{G}_{d}:=\big{[}\ket{+}\!\!\bra{+}\oplus 0,M\big{)}=\big{\{}(1-\lambda)\ket{+}\!\!\bra{+}\oplus\lambda\colon\lambda\in[0,\frac{1}{2})\big{\}}$
and continuous at every point in the complement. (Unitary similarity with respect to $U\oplus 1$ matches $\mathcal{R}$ with the problem considered in [52], as the similarity with respect to $U:=\exp(\mathrm{i}\frac{\pi}{3\sqrt{3}}(X+Y+Z))$ permutes the Pauli matrices $X,Y,Z$ cyclicly.) Lemma 9.7 then shows that the restricted rI-projection $\Pi|_{\mathcal{D}(\mathcal{A})}$ is discontinuous at every point in $\mathcal{G}_{d}$ and continuous at every point in the complement. Now Rem. 9.8 proves that the rI-projection $\Pi$ is discontinuous at every point in $\pi_{\mathcal{A}}^{-1}(\mathcal{G}_{d})$ and continuous at every point in the complement. A second application of Lemma 9.7 finally shows that the entropy distance $d$ is discontinuous at every point in $\pi_{\mathcal{A}}^{-1}(\mathcal{G}_{d})$ and continuous at every point in the complement.
The points in the set $\pi_{\mathcal{A}}^{-1}(\mathcal{G}_{d})$ are explicitly described in Ex. 8.3 b). The result is consistent with the inference $\Psi$ being discontinuous at $M$ and continuous at every point different from $M$ (see Ex. 9.3 c)) as per Lemma 9.7 b).
Acknowledgments. The author is grateful to several people for discussions related to this work, among them are Achim Clausing, Chi-Kwong Li, Tim Netzer, Ivan Todorov, and Ilya M. Spitkovsky. The author gave a lecture on the results of this paper in the special session “Numerical Ranges” of the conference IWOTA 2024 in Canterbury, UK, and wishes to thank the organizers for this opportunity.

### v2

@@
 Key words and phrases:
-open map, numerical range, matrix system, quantum state
+quantum state, matrix system, open map, numerical range
 2020 Mathematics Subject Classification:
-46A30, 15A60, 47A12, 81P16
+Primary 81P16, 94A15, 46A30; Secondary 47A12, 15A60
 In honor of Ilya Matveevich Spitkovsky, for his 70th birthday.
@@
 $\omega_{\lambda}:=r_{\mathrm{M}_{3}}\big{[}(1-\lambda)\ket{+}\!\!\bra{+}\oplus\lambda\big{]}\,,\quad\lambda\in(0,1]\,.$
-So $\mathcal{O}|_{\mathcal{R}}:=\{\ell|_{\mathcal{R}}:\ell\in\mathcal{O}\}$ contains $\omega_{\lambda}|_{\mathcal{R}}$ but none of the restriction $\ell_{\theta}|_{\mathcal{R}}$ of
+So $\mathcal{O}|_{\mathcal{R}}:=\{\ell|_{\mathcal{R}}:\ell\in\mathcal{O}\}$ contains $\omega_{\lambda}|_{\mathcal{R}}$ but none of the restrictions $\ell_{\theta}|_{\mathcal{R}}$ of
 $\ell_{\theta}:=r_{\mathrm{M}_{3}}\big{[}\frac{1}{2}\big{(}\mathds{1}_{2}+\cos(\theta)X+\sin(\theta)Z\big{)}\oplus 0\big{]}\,,\quad\theta\in(0,2\pi)\,.$
@@
 [29, 30] first studied a problem of numerical ranges closely related to (1.1).
-The problem (1.1) was studied by Weis [52, 53] and Rodman et al.
+The problem (1.1) itself was studied by Weis [53, 54] and Rodman et al.
 [46] when $\mathcal{D}(\mathcal{R})$ is replaced with the affinely isomorphic joint numerical range.
@@
 Thm.
-1.2 can simplify the problem (1.1), and related continuity problems, if $\mathcal{R}$ is included in a *-subalgebra of $\mathrm{M}_{n}$ smaller than $\mathrm{M}_{n}$, as we show in Sec.
+1.2 can simplify the problem (1.1), and related continuity problems, if $\mathcal{R}$ is included in a *-subalgebra of $\mathrm{M}_{n}$ smaller than $\mathrm{M}_{n}$ as we show in Sec.
 8.
@@
 In the sequel, we refer to matrix systems and *-subalgebras synonymously as complex matrix systems and complex *-subalgebras, respectively.
-We introduce their real counterparts because real *-subalgebras have a greater variety of state spaces than the complex ones.
+In addition, we introduce real *-subalgebras as they have a greater variety of state spaces than the complex ones.
+See, however, Rem.
+2.4.
 A real matrix system on $\mathbb{C}^{n}$ is a real linear subspace $\mathcal{R}$ of $\mathrm{M}_{n}$ that is self-adjoint and contains $\mathds{1}_{n}$.
@@
 Conversely, any real matrix system $\mathcal{R}$ included in $\mathcal{H}(\mathrm{M}_{n})$ has the same same set of hermitian matrices and the same set of density matrices as the complex matrix system $\mathcal{R}\oplus\mathrm{i}\mathcal{R}$.
-The set of density matrices of a real or complex matrix system $\mathcal{R}$ is affinely isomorphic the state space of $\mathcal{R}$ by Lemma 2.1 or 2.2, respectively.
+The set of density matrices of a real or complex matrix system $\mathcal{R}$ is affinely isomorphic to the state space of $\mathcal{R}$ by Lemma 2.1 or 2.2, respectively.
 Let $\mathcal{R}_{1},\mathcal{R}_{2}$ be real matrix systems on $\mathbb{C}^{n}$ and let $\mathcal{R}_{2}\subset\mathcal{R}_{1}$.
@@
 III.2.1] as the space of hermitian matrices $\mathcal{H}_{2}$ is a Euclidean Jordan algebra with Jordan product $A\circ B=\frac{1}{2}(AB+BA)$ and inner product $(A,B)\mapsto\mathrm{Re}\langle A,B\rangle$, $A,B\in\mathcal{H}_{2}$.
-So, $\mathcal{C}_{2}^{\vee}=\mathcal{C}_{2}\subset\mathcal{C}_{1}\subset\mathcal{C}_{1}^{\vee}$ proves $\mathcal{C}_{2}^{\vee}=\mathcal{C}_{2}$, which implies (2.4).
+So, $\mathcal{C}_{2}^{\vee}=\mathcal{C}_{2}\subset\mathcal{C}_{1}\subset\mathcal{C}_{1}^{\vee}$ proves $\mathcal{D}_{2}\subset\mathcal{D}_{1}$, which implies (2.4).
 See also the Notes to Chapter 6 in [2].
@@
 Proof.
-As $K:=K_{1}\oplus_{\mathrm{c}}\cdots\oplus_{\mathrm{c}}K_{m}$ is a metric space, it suffices to show that there are arbitrary small family members and $O:=O_{I}(\epsilon,(A_{i})_{i\in I})$ is open.
+As $K:=K_{1}\oplus_{\mathrm{c}}\cdots\oplus_{\mathrm{c}}K_{m}$ is a metric space, it suffices to show that $O:=O_{I}(\epsilon,(A_{i})_{i\in I})$ is open and there are arbitrary small sets of this form.
 Let $B_{i}:=A_{i}$ if $i\in I$ and $B_{i}:=K_{i}$ if $i\not\in I$, $i=1,\dots,m$. Then the set
@@
 It is important to observe that $\frac{1}{2}(\mathds{1}_{2}+\cos(\theta)X+\sin(\theta)Z)\oplus 0$ and $\cos(\theta)(X\oplus 1)+\sin(\theta)Z\oplus 0$ are matrices in $\mathrm{M}_{2}\oplus\mathrm{M}_{1}$ for all $\theta\in\mathbb{R}$.
-The map $\pi$ is open at every point in the complement of $\mathcal{G}_{0}$ because $\mathcal{D}(\mathrm{M}_{2}\oplus\mathrm{M}_{1})$ is a cone over a ball, see [52, Lemma 4.17] for a detailed proof.
+The map $\pi$ is open at every point in the complement of $\mathcal{G}_{0}$ because $\mathcal{D}(\mathrm{M}_{2}\oplus\mathrm{M}_{1})$ is a cone over a ball, see [53, Lemma 4.17] for a detailed proof.
   2. b)
@@
 2.9, the value $(1/3,0,\sqrt{8/9})$ confirms that $(\mathds{1}_{3}+\lambda\,Z\oplus 0)/3$ is contained in $\mathcal{D}(\mathcal{R})$ if and only if $|\lambda|\leq\sqrt{2}$.
+It is instructive to look at Ex. 3.5 from the perspective of a real *-subalgebra of $\mathrm{M}_{3}$, whose state space can be visualized in three-space.
 Example 3.6.
@@
 over $M$.
-The map $\pi$ is not open at any point in the half-open segment $\mathcal{G}_{0}$ and open at every point in the complement.
+The map $\pi$ is not open at any point in $\mathcal{G}_{0}=\mathcal{G}\setminus\{\ket{+}\!\!\bra{+}\oplus 0\}$ and open at every point in the complement of $\mathcal{G}_{0}$.
 The lack of openness can be visualized graphically in three-space by observing that $\pi$ projects the cone $\mathcal{D}(\mathrm{M}_{2}(\mathbb{R})\oplus\mathrm{M}_{1}(\mathbb{R}))$ along its generatrix $\mathcal{G}$ to the ellipse $\mathcal{D}(\mathcal{R})$.
@@
 We use the same label $f_{A}$ to denote the map
-$f_{A}:\mathbb{C}S^{n}\to W(A)\,,\quad\ket{\varphi}\mapsto\braket{\varphi}{A\varphi}\,.$ (4.2)
+$f_{A}:\mathbb{C}S^{n}\to W(A)\,,\quad\ket{\varphi}\mapsto\braket{\varphi}{A\varphi}\,.$
 Minkowski’s theorem [47] asserts that every compact convex set is the convex hull of its extreme points.
@@
 Every extreme point of $W(A)$ is contained in $\operatorname{Img}(z_{i})$ for some $i=1,\ldots,n$.
-If $z\in W(A)$ is not an extreme point of $W(A)$, then $f_{A}^{-1}$ is strongly continuous at $z$ [18, Thm.
+If $z\in W(A)$ is not an extreme point of $W(A)$, then $f_{A}^{-1}$ is strongly continuous at $z$ [17, Thm.
 4].
@@
 Strong continuity is connected to the openness of a linear map.
-For all $z\in W(A)$, the multi-valued map $f_{A}^{-1}$ is strongly continuous at $z$ if and only if the restricted linear map $v:\mathcal{D}(\mathrm{M}_{n})\to V(A_{1},A_{2})$, introduced in (4.1), is open at every point in the fiber $v^{-1}[(\mathrm{Re}(z),\mathrm{Im}(z)){}^{\mathrm{T}}]$, see [53, Coro.
+For all $z\in W(A)$, the multi-valued map $f_{A}^{-1}$ is strongly continuous at $z$ if and only if the restricted linear map $v:\mathcal{D}(\mathrm{M}_{n})\to V(A_{1},A_{2})$, introduced in (4.1), is open at every point in the fiber $v^{-1}[(\mathrm{Re}(z),\mathrm{Im}(z)){}^{\mathrm{T}}]$, see [54, Coro.
 5.2].
@@
     The stability problem of a finite-dimensional compact convex set $K$ is completely solved [40].
-The $d$-skeleton of $K$ is the union of all faces [^3] A face [45] of a convex set $K$ is a convex subset $F\subset K$ such that $x,y\in F$ is implied by $(1-s)x+sy\in F$ for all $x,y\in K$ and $s\in(0,1)$.
+The $d$-skeleton of $K$ is defined as the union of all faces [^3] A face [45] of a convex set $K$ is a convex subset $F\subset K$ such that $x,y\in F$ is implied by $(1-s)x+sy\in F$ for all $x,y\in K$ and $s\in(0,1)$.
 of $K$ of dimension at most $d$.
@@
 The closedness follows from three arguments: First, the set $\mathcal{D}(\mathrm{M}_{n})$ is a compact convex set of dimension $n^{2}-1$.
-Second, every nonempty face of $\mathrm{M}_{n}$ is unitarily similar to $\mathcal{D}(\mathrm{M}_{l})\oplus\{0\}$ for some positive integer $l$, and third, and the unitary group $U(n)$ is compact.
+Second, every nonempty face of $\mathrm{M}_{n}$ is unitarily similar to $\mathcal{D}(\mathrm{M}_{l})\oplus\{0\}$ for some positive integer $l$, and third, the unitary group $U(n)$ is compact.
   3. c)
@@
 If $K$ is a stable convex set, then the arithmetic mean map
-$K^{\times k}\to K\,,\quad(x_{1},\dots,x_{k})\mapsto\frac{1}{k}\sum_{i=1}^{k}x_{i}$
-the $k$-fold cartesian product of $K$ is open for all positive integers $k$. More generally, for every tuple $(s_{1},\dots,s_{k})$ of nonnegative real numbers adding up to one, the map
+$K^{\times k}\to K\,,\quad(x_{1},\dots,x_{k})\mapsto\frac{1}{k}\sum_{i=1}^{k}x_{i}\,,$
+defined on the $k$-fold cartesian product of $K$, is open for all positive integers $k$. More generally, for every tuple $(s_{1},\dots,s_{k})$ of nonnegative real numbers adding up to one, the map
 $K^{\times k}\to K\,,\quad(x_{1},\dots,x_{k})\mapsto\sum_{i=1}^{k}s_{i}x_{i}$
@@
 For every $\rho\in\mathcal{D}(\mathrm{M}_{n})$ the map $\pi$ is open at $\rho$ if and only if $\pi_{\mathcal{R}}$ is open at $\pi_{\mathcal{A}}(\rho)$.
-As a word of explication, the map $\pi$ factors by Lemma 2.5.
-The claim follows from Lemma 8.1 a), by letting $\mathcal{R}_{1}:=\mathrm{M}_{n}$ and $\mathcal{R}_{2}:=\mathcal{A}$, and by taking $\pi_{\mathcal{R}}$ as the map $f:\mathcal{D}(\mathcal{R}_{2})\to T$.
+Indeed, the map $\pi$ factors by Lemma 2.5.
+The second claim follows from Lemma 8.1 a), by letting $\mathcal{R}_{1}:=\mathrm{M}_{n}$ and $\mathcal{R}_{2}:=\mathcal{A}$, and by taking $\pi_{\mathcal{R}}$ as the map $f:\mathcal{D}(\mathcal{R}_{2})\to T$.
 The assumptions of the lemma are met since $\pi_{\mathcal{A}}$ is open by Thm.
 1.2.
-The following two examples are concerned with openness. Continuity is a topic of Sec. 9 below.
+The following examples demonstrate the use of Rem. 8.2. Continuity problems are a topic of Sec. 9 below.
 Example 8.3.
-Let $\mathcal{R}:=\operatorname{span}_{\mathbb{R}}(\mathds{1}_{3},X\oplus 1,Z\oplus 0)$.
+Keeping the notation of $\pi=\pi_{\mathcal{R}}\circ\pi_{\mathcal{A}}$ from Rem.
+8.2, we take $\mathcal{R}:=\operatorname{span}_{\mathbb{R}}(\mathds{1}_{3},X\oplus 1,Z\oplus 0)$ and $\mathcal{A}:=\mathrm{M}_{2}\oplus\mathrm{M}_{1}$.
   1. a)
-    Rem.
-8.2 solves the openness problem of the orthogonal projection $\pi:\mathcal{D}(\mathrm{M}_{3})\to\mathcal{D}(\mathcal{R})$ to $\mathcal{R}$, based on the solution of the openness problem of $\pi_{\mathcal{R}}:\mathcal{D}(\mathcal{A})\to\mathcal{D}(\mathcal{R})$, where $\mathcal{A}:=\mathrm{M}_{2}\oplus\mathrm{M}_{1}$.
     By Ex.
@@
 6.2), one verifies that the fiber of $\pi_{\mathcal{A}}$ over this point is the set of all matrices
-    $\rho(\lambda,z):=\begin{pmatrix}(1-\lambda)\ket{+}\!\!\bra{+}&z\ket{+}\\ \overline{z}\bra{+}&\lambda\end{pmatrix}\,,\quad z\in\mathbb{C}\,,\quad|z|^{2}\leq\lambda(1-\lambda)\,,$ (8.1)
+    $\rho(\lambda,z):=\begin{pmatrix}(1-\lambda)\ket{+}\!\!\bra{+}&z\ket{+}\\ \overline{z}\bra{+}&\lambda\end{pmatrix}\,,\quad z\in\mathbb{C}\,,\quad|z|^{2}\leq\lambda(1-\lambda)\,,$
     where $|z|$ denotes the absolute value of $z\in\mathbb{C}$. In conclusion, the orthogonal projection $\pi:\mathcal{D}(\mathrm{M}_{3})\to\mathcal{D}(\mathcal{R})$ is not open at any point of
@@
 The preferred posterior is that which minimizes the relative entropy from the prior subject to the available constraints, see Chap.
-8 in Caticha’s book [13], and the references therein.
-Whereas the maximum entropy method has a sound axiomatic basis, the analogous inference method for quantum states lacks an axiomatic foundation and remains a controversial topic [3].
-Analytic properties of quantum mechanical inference are interesting despite unsolved foundational issues.
-Notably, quantum inference maps have discontinuities [55] that aroused interest in theoretical physics [14, 26, 36, 56].
+8 in Caticha’s book [12] and the references therein.
+An analogous quantum mechanical inference method can be defined by replacing probability distributions with density matrices and the standard relative entropy with the Umegaki relative entropy.
+The axiomatic foundations of the maximum entropy inference method were settled for probability distributions in the 1980’s, see Chap.
+6 in [12, pp.
+157–160].
+More than 30 years later, the search for axioms of the quantum inference [3] found a convincing answer in the work of Vanslette [51].
 Linear constraints [^4] Linear constraints can be defined in terms of expectation values.
@@
 The relative entropy $S:\mathcal{D}(\mathrm{M}_{n})\times\mathcal{D}(\mathrm{M}_{n})\to[0,+\infty]$ is an asymmetric distance.
-It is defined as $S(\rho_{1},\rho_{2}):=\operatorname{tr}[\rho_{1}(\log(\rho_{1})-\log(\rho_{2}))]$ if the range of $\rho_{1}$ is included in the range of $\rho_{2}$ and as $S(\rho_{1},\rho_{2}):=+\infty$ otherwise, for all $\rho_{1},\rho_{2}\in\mathcal{D}(\mathrm{M}_{n})$.
+It is defined by $S(\rho_{1},\rho_{2}):=\operatorname{tr}[\rho_{1}(\log(\rho_{1})-\log(\rho_{2}))]$ if the range of $\rho_{1}$ is included in the range of $\rho_{2}$ and by $S(\rho_{1},\rho_{2}):=+\infty$ otherwise, for all $\rho_{1},\rho_{2}\in\mathcal{D}(\mathrm{M}_{n})$.
 Let $\tau\in\mathcal{D}(\mathrm{M}_{n})$, the prior, be a density matrix of maximal rank $n$.
@@
 $\Psi_{\tau}:\mathcal{D}(\mathcal{R})\to\mathcal{D}(\mathrm{M}_{n})\,,\quad\sigma\mapsto\operatorname{argmax}_{\rho\in\pi^{-1}(\sigma)}\phi_{\tau}(\rho)$ (9.1)
-is well defined, see [52, Def. 1.1] and the references therein.
-We quote [52, Thm. 4.9] under the assumptions of Ex. 9.1.
+is well defined, see [53, Def.
+1.1] and the references therein.
+Discontinuities of this inference map [56] aroused interest in theoretical physics [13, 26, 36, 57].
+The map (9.1) is continuous for commutative real matrix systems $\mathcal{R}$, for example for the inference of probability distributions mentioned above.
+We quote [53, Thm. 4.9].
 Theorem 9.2.
-For any $\sigma\in\mathcal{D}(\mathcal{R})$, the inference map $\Psi_{\tau}:\mathcal{D}(\mathcal{R})\to\mathcal{D}(\mathrm{M}_{n})$ is continuous at $\sigma$ if and only if the orthogonal projection $\pi:\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathcal{R})$ is open at $\Psi_{\tau}(\sigma)$.
+Let $\sigma\in\mathcal{D}(\mathcal{R})$. Then $\Psi_{\tau}$ is continuous at $\sigma$ if and only if $\pi$ is open at $\Psi_{\tau}(\sigma)$.
+In what follows, continuity and openness problems will be simplified by factoring $\pi=\pi_{\mathcal{R}}\circ\pi_{\mathcal{A}}$ through a *-subalgebra $\mathcal{A}$ of $\mathrm{M}_{n}$ that includes $\mathcal{R}$, using the notation of Rem.
+8.2.
 Example 9.3 (Maximum entropy inference II).
-We resume Ex.
-9.1.
-Thm.
-9.2 translates the continuity problem of the inference map $\Psi_{\tau}$ into an openness problem of the orthogonal projection $\pi$.
-Whether Rem.
-8.2 can simplify this openness problem depends on the location of the posteriors.
-The range of $\Psi_{\tau}$ can be described explicitly. The manifold
-$\mathcal{E}_{\tau}(\mathcal{R}):=\big{\{}\frac{e^{\log(\tau)+A}}{\operatorname{tr}(e^{\log(\tau)+A})}\mid A\in\mathcal{R}\big{\}}$ (9.2)
-is known as a Gibbsian family or exponential family, see [55, 38, 52, 43] and the references therein (and recall $\mathcal{R}=\mathcal{H}(\mathcal{R})$).
-The manifold $\mathcal{E}_{\tau}(\mathcal{R})$ is included in $\Psi_{\tau}(\mathcal{D}(\mathcal{R}))$.
-But not just that, the set $\Psi_{\tau}(\mathcal{D}(\mathcal{R}))$ of maximum entropy inference states equals
-$\{\rho_{1}\in\mathcal{D}(\mathrm{M}_{n}):\inf_{\rho_{2}\in\mathcal{E}_{\tau}(\mathcal{R})}S(\rho_{1},\rho_{2})=0\}\,,$ (9.3)
-which is a subset of the Euclidean closure of $\mathcal{E}_{\tau}(\mathcal{R})$, see (D5) in [52].
-The set (9.3) is called reverse information closure or the rI-closure [19] of $\mathcal{E}(\mathcal{R})$.
-Let $\mathcal{A}$ denote the smallest *-subalgebra $\mathcal{A}$ of $\mathrm{M}_{n}$ that contains $\mathcal{R}$.
+Before simplifying the openness problem of $\pi$, we describe the set $\Psi_{\tau}(\mathcal{D}(\mathcal{R}))$ of posteriors, as the openness only matters for points in this set by Thm.
+9.2.
+Recalling $\mathcal{R}=\mathcal{H}(\mathcal{R})$, we define
+$\mathcal{E}_{\tau}(\mathcal{R}):=\big{\{}\frac{e^{\log(\tau)+A}}{\operatorname{tr}(e^{\log(\tau)+A})}\mid A\in\mathcal{R}\big{\}}\,.$
+The manifold $\mathcal{E}_{\tau}(\mathcal{R})$ is known as a Gibbsian family or exponential family, see [56, 38, 53, 43] and the references therein. By (D5) in [53] we have
+$\Psi_{\tau}(\mathcal{D}(\mathcal{R}))=\{\rho_{1}\in\mathcal{D}(\mathrm{M}_{n}):\inf_{\rho_{2}\in\mathcal{E}_{\tau}(\mathcal{R})}S(\rho_{1},\rho_{2})=0\}\,.$ (9.2)
+The right-hand side of (9.2) is the reverse information closure or rI-closure [18] of $\mathcal{E}_{\tau}(\mathcal{R})$, which is a subset of the Euclidean closure of $\mathcal{E}_{\tau}(\mathcal{R})$.
   1. a)
-    Rem. 8.2 is of no avail if $\tau=\mathds{1}_{n}/n$ is the uniform prior. In this case,
+    If $\tau=\mathds{1}_{n}/n$ is the uniform prior, then
     $\phi_{\tau}(\rho)=-S(\rho,\mathds{1}_{n}/n)=S(\rho)-\log(n)\,,\quad\rho\in\mathcal{D}(\mathrm{M}_{n})$
     is the von Neumann entropy $S(\rho):=-\operatorname{tr}[\rho\log(\rho)]$ up to a constant.
-By functional calculus, the exponential family $\mathcal{E}_{\tau}(\mathcal{R})$ is included in $\mathcal{A}$ and so is the set of inference states $\Psi_{\tau}(\mathcal{D}(\mathcal{R}))$ as per the formula (9.3), because $\mathcal{A}$ is closed.
+By functional calculus, $\mathcal{E}_{\tau}(\mathcal{R})$ is included in $\mathcal{A}$ and so is the set of posteriors $\Psi_{\tau}(\mathcal{D}(\mathcal{R}))$ as per (9.2), because $\mathcal{A}$ is closed.
 Rem.
-8.2 is of no help since all posteriors lie in $\mathcal{A}$ anyways.
-The same conclusion is true for every prior $\tau$ in $\mathcal{A}$ replacing $\mathds{1}_{n}/n$.
+8.2 then shows that for every $\rho\in\Psi_{\tau}(\mathcal{D}(\mathcal{R}))$ the openness of $\pi$ at $\rho$ is equivalent to the openness of $\pi_{\mathcal{R}}=\pi|_{\mathcal{D}(\mathcal{A})}$ at $\rho$ (the same conclusion is true for every prior $\tau$ in $\mathcal{A}$).
+This can simplify the problem if $\mathcal{A}$ has a simpler structure or a smaller dimension than $\mathrm{M}_{n}$.
+An example is given in Ex.
+8.4 above.
   2. b)
-    If the prior $\tau$ lies outside of $\mathcal{A}$ then the exponential family $\mathcal{E}_{\tau}(\mathcal{R})$ is disjoint from $\mathcal{D}(\mathcal{A})$, again by functional calculus.
+    If the prior $\tau$ lies outside of $\mathcal{A}$ then $\mathcal{E}_{\tau}(\mathcal{R})$ is disjoint from $\mathcal{D}(\mathcal{A})$, again by functional calculus.
 Rem.
-8.2 brings an advantage if the orthogonal projection $\pi_{\mathcal{R}}:\mathcal{D}(\mathcal{A})\to\mathcal{D}(\mathcal{R})$ is easier to analyze than $\pi:\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathcal{R})$.
-An example is given in Ex.
-8.4 where $\mathcal{A}$ has a simple structure and a much smaller dimension than $\mathrm{M}_{n}$.
+8.2 then shows that for every $\rho\in\Psi_{\tau}(\mathcal{D}(\mathcal{R}))$ the openness of $\pi$ at $\rho$ is equivalent to the openness of $\pi_{\mathcal{R}}$ at $\pi_{\mathcal{A}}(\rho)$.
+This is an even greater simplification than in part a) above, because the analysis of $\pi$ on $\mathcal{D}(\mathrm{M}_{n})$ is reduced to that of $\pi_{\mathcal{R}}$ on $\mathcal{D}(\mathcal{A})$.
   3. c)
-    The real matrix system $\mathcal{R}:=\operatorname{span}_{\mathbb{R}}(\mathds{1}_{3},X\oplus 1,Z\oplus 0)$ in Ex.
-8.3 has the property that the orthogonal projection $\pi:\mathcal{D}(\mathrm{M}_{3})\to\mathcal{D}(\mathcal{R})$ is not open anywhere in the relative interior of the fiber over $M:=\frac{1}{2}(\ket{+}\!\!\bra{+}\oplus 1)$ and open at every point in the complement of that fiber.
-For every $\sigma\in\mathcal{D}(\mathcal{R})$ and prior $\tau$, the posterior $\Psi_{\tau}(\sigma)$ is contained in the relative interior of the fiber $\pi^{-1}(\sigma)$ over every $\sigma$ by Coro.
+    Independently of $\mathcal{A}$, it sometimes helps that the posterior $\Psi_{\tau}(\sigma)$ is contained in the relative interior of the fiber $\pi^{-1}(\sigma)$ over $\sigma$ for every $\sigma\in\mathcal{D}(\mathcal{R})$ and prior $\tau$ by Coro.
 5.7 and Lemma 5.8 in [53].
+As an example, the orthogonal projection $\pi:\mathcal{D}(\mathrm{M}_{3})\to\mathcal{D}(\mathcal{R})$ to the real matrix system $\mathcal{R}$ in Ex.
+8.3 is not open anywhere in the relative interior of the fiber over a certain point $M$ and open at every point in the complement of that fiber.
 Thm.
-9.2 then shows that $\Psi_{\tau}$ is discontinuous at $M$ and continuous at every point different from $M$ (for all priors $\tau$).
+9.2 then shows that $\Psi_{\tau}$ is discontinuous at $M$ and continuous everywhere else in the ellipse $\mathcal{D}(\mathcal{R})$ for every prior $\tau$.
+From here on, we assume the prior $\tau:=\mathds{1}_{n}/n$ be uniform. Using the von Neumann entropy $S=\phi_{\tau}+\log(n)$, we write the inference map (9.1) from Ex. 9.1 as
+$\Psi:\mathcal{D}(\mathcal{R})\to\mathcal{D}(\mathrm{M}_{n})\,,\quad\sigma\mapsto\operatorname{argmax}_{\rho\in\pi^{-1}(\sigma)}S(\rho)\,.$
+We also write $\mathcal{E}(\mathcal{R}):=\mathcal{E}_{\tau}(\mathcal{R})$ for the exponential family of Ex. 9.3.
 Example 9.4 (Maximum entropy inference III).
@@
 A $\mathfrak{g}$-local Hamiltonian is a hermitian matrix in $\mathcal{A}_{\Omega}$ of the form
-$\sum_{\nu\in\mathfrak{g}}A_{\nu}\otimes\mathds{1}_{\bar{\nu}}\,,\quad A_{\nu}\in\mathcal{H}(\mathcal{A}_{\nu})\,,\quad\nu\in\mathfrak{g}\,.$ (9.4)
+$\sum_{\nu\in\mathfrak{g}}A_{\nu}\otimes\mathds{1}_{\bar{\nu}}\,,\quad A_{\nu}\in\mathcal{H}(\mathcal{A}_{\nu})\,,\quad\nu\in\mathfrak{g}\,.$
 We denote the real matrix system of all $\mathfrak{g}$-local Hamiltonians by $\mathcal{R}_{\mathfrak{g}}$ and the orthogonal projection by $\pi_{\mathfrak{g}}:\mathcal{D}(\mathcal{A}_{\Omega})\to\mathcal{D}(\mathcal{R}_{\mathfrak{g}})$.
@@
 Let
-$\operatorname{red}_{\mathfrak{g}}:\mathcal{A}_{\Omega}\to\prod_{\nu\in\mathfrak{g}}\mathcal{A}_{\nu}\,,\quad A\mapsto[\operatorname{Tr}_{\bar{\nu}}(A)]_{\nu\in\mathfrak{g}}$ (9.5)
-denote the map from $\mathcal{A}_{\Omega}$ to the cartesian product of the algebras $(\mathcal{A}_{\nu})_{\nu\in\mathfrak{g}}$ that assigns partial traces to the subsystems specified by the members of $\mathfrak{g}$.
-Linear constraints on $\mathcal{D}(\mathcal{A}_{\Omega})$ have been defined in terms of reduced density matrices, see [38, 14] and [58, Sec.
+$\operatorname{red}_{\mathfrak{g}}:\mathcal{A}_{\Omega}\to\prod_{\nu\in\mathfrak{g}}\mathcal{A}_{\nu}\,,\quad A\mapsto[\operatorname{tr}_{\bar{\nu}}(A)]_{\nu\in\mathfrak{g}}$
+denote the map from $\mathcal{A}_{\Omega}$ to the cartesian product of the algebras $(\mathcal{A}_{\nu})_{\nu\in\mathfrak{g}}$ that assigns reduced density matrices.
+Linear constraints on $\mathcal{D}(\mathcal{A}_{\Omega})$ have been defined in terms of reduced density matrices, see [38, 13] and [59, Sec.
 1.4.2].
 This is formalized in the following diagram, which commutes by formula (19) in [55].
-(Obvious restrictions of $\operatorname{red}_{\mathfrak{g}}$ are omitted.)
+(Obvious restrictions of the domain and codomain of $\operatorname{red}_{\mathfrak{g}}$ are omitted in the sequel.)
 ${\mathcal{D}(\mathcal{A}_{\Omega})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{\mathfrak{g}}}$$\scriptstyle{\operatorname{red}_{\mathfrak{g}}}$${\operatorname{red}_{\mathfrak{g}}[\mathcal{D}(\mathcal{A}_{\Omega})]}$${\mathcal{D}(\mathcal{R}_{\mathfrak{g}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\operatorname{red}_{\mathfrak{g}}}$
-The commutative diagram shows that the fiber $\pi_{\mathfrak{g}}^{-1}(\sigma)$ over $\sigma\in\mathcal{D}(\mathcal{R}_{\mathfrak{g}})$ is the set of all $\rho\in\mathcal{D}(\mathcal{A}_{\Omega})$ such that $\operatorname{red}_{\mathfrak{g}}(\rho)=\operatorname{red}_{\mathfrak{g}}(\sigma)$.
+The fiber $\pi_{\mathfrak{g}}^{-1}(\sigma)$ over $\sigma\in\mathcal{D}(\mathcal{R}_{\mathfrak{g}})$ is the set of all $\rho\in\mathcal{D}(\mathcal{A}_{\Omega})$ such that $\operatorname{red}_{\mathfrak{g}}(\rho)=\operatorname{red}_{\mathfrak{g}}(\sigma)$.
 Thm.
-9.2 proves that the pullback $\Xi_{\tau}:=\Psi_{\tau}\circ\operatorname{red}_{\mathfrak{g}}^{-1}$ of the inference map $\Psi_{\tau}$ under $\operatorname{red}_{\mathfrak{g}}^{-1}$ is continuous at $(\rho_{\nu})_{\nu\in\mathfrak{g}}\in\operatorname{red}_{\mathfrak{g}}[\mathcal{D}(\mathcal{A}_{\Omega})]$ if and only if $\operatorname{red}_{\mathfrak{g}}$ is open at $\Xi_{\tau}[(\rho_{\nu})_{\nu\in\mathfrak{g}}]\in\mathcal{D}(\mathcal{A}_{\Omega})$.
-A discontinuity of $\Xi_{\tau}$ (with uniform prior $\tau$) was reported for $N=3$ qubits and $\mathfrak{g}:=\{\{1,2\},\{2,3\},\{3,1\}\}$ at $(\rho_{\nu})_{\nu\in\mathfrak{g}}\in\operatorname{red}_{\mathfrak{g}}[\mathcal{D}(\mathcal{A}_{\Omega})]$, where
-$\rho_{\nu}:=\frac{1}{2}(\ket{00}\!\!\bra{00}+\ket{11}\!\!\bra{11})\,,\quad\nu\in\mathfrak{g}\,.$
-This result by Chen et al.
-[14, Ex.
-4] has an interesting interpretation.
-The map $\operatorname{red}_{\mathfrak{g}}$ being open [^5] The openness of $\operatorname{red}_{\mathfrak{g}}$ at $\rho\in\mathcal{D}(\mathcal{A}_{\Omega})$ is a priori weaker than the continuity of $\Xi_{\tau}$ at $\operatorname{red}_{\mathfrak{g}}(\rho)$.
-The continuity means that any sufficiently small change of $\operatorname{red}_{\mathfrak{g}}(\rho)$ can be matched by an arbitrarily small change of $\rho$ inside the range of $\Xi_{\tau}$ and not just anywhere within $\mathcal{D}(\mathcal{A}_{\Omega})$.
+9.2 proves that the pullback $\Xi:=\Psi\circ\operatorname{red}_{\mathfrak{g}}^{-1}$ of the inference map $\Psi$ under $\operatorname{red}_{\mathfrak{g}}^{-1}$ is continuous at $(\rho_{\nu})_{\nu\in\mathfrak{g}}\in\operatorname{red}_{\mathfrak{g}}[\mathcal{D}(\mathcal{A}_{\Omega})]$ if and only if $\operatorname{red}_{\mathfrak{g}}$ is open at $\Xi[(\rho_{\nu})_{\nu\in\mathfrak{g}}]\in\mathcal{D}(\mathcal{A}_{\Omega})$.
+Chen et al.
+[13, Ex.
+4] discovered a discontinuity of $\Xi$ for $N=3$ qubits and $\mathfrak{g}:=\{\{1,2\},\{2,3\},\{3,1\}\}$ at $(\rho_{\nu})_{\nu\in\mathfrak{g}}\in\operatorname{red}_{\mathfrak{g}}[\mathcal{D}(\mathcal{A}_{\Omega})]$, where
+$\rho_{\nu}:=\frac{1}{2}(\ket{00}\!\!\bra{00}+\ket{11}\!\!\bra{11})\,,\quad\nu\in\mathfrak{g}$
+and they offered an interesting interpretation in terms of phase transitions.
+The map $\operatorname{red}_{\mathfrak{g}}$ being open [^5] The openness of $\operatorname{red}_{\mathfrak{g}}$ at $\rho\in\mathcal{D}(\mathcal{A}_{\Omega})$ is a priori weaker than the continuity of $\Xi$ at $\operatorname{red}_{\mathfrak{g}}(\rho)$.
+The continuity means that any sufficiently small change of $\operatorname{red}_{\mathfrak{g}}(\rho)$ can be matched by an arbitrarily small change of $\rho$ inside the image of $\Xi$ and not just anywhere in $\mathcal{D}(\mathcal{A}_{\Omega})$.
 Somewhat surprisingly, the two propositions are equivalent by Thm.
 9.2.
-at $\rho\in\mathcal{D}(\mathcal{A}_{\Omega})$ means that any sufficiently small change of $\operatorname{red}_{\mathfrak{g}}(\rho)$ within $\operatorname{red}_{\mathfrak{g}}[\mathcal{D}(\mathcal{A}_{\Omega})]$ is matched by an arbitrarily small change of $\rho$ within $\mathcal{D}(\mathcal{A}_{\Omega})$.
-Conversely, if the openness fails, then there are arbitrarily small changes of $\operatorname{red}_{\mathfrak{g}}(\rho)$ that are only matched by changes of $\rho$ beyond some positive threshold (in the metric sense).
+at $\rho\in\mathcal{D}(\mathcal{A}_{\Omega})$ means that any sufficiently small change of $\operatorname{red}_{\mathfrak{g}}(\rho)$ in $\operatorname{red}_{\mathfrak{g}}[\mathcal{D}(\mathcal{A}_{\Omega})]$ is matched by an arbitrarily small change of $\rho$ within $\mathcal{D}(\mathcal{A}_{\Omega})$.
+Conversely, if the openness fails, then there are arbitrarily small changes of $\operatorname{red}_{\mathfrak{g}}(\rho)$ that can only be matched by changes of $\rho$ beyond some strictly positive threshold (in the metric sense).
 Loosely speaking, a small change of a subsystem abruptly changes the entire system.
-Such behavior is associated with phase transitions [14].
-It motivates every attempt to study the openness problem of $\pi_{\mathfrak{g}}$.
+Such behavior is associated with phase transitions.
+It motivates every attempt to study the openness of $\pi_{\mathfrak{g}}$.
 This should be done by tuning the interaction pattern $\mathfrak{g}$ to a concrete system.
-Whether enclosing $\mathcal{R}_{\mathfrak{g}}$ into a *-subalgebra of $\mathcal{A}_{\Omega}$ could simplify the problem, as suggested in Rem.
-8.2, is not yet clarified.
-Focusing on the continuity of correlations measures, we begin with theory.
+Whether enclosing $\mathcal{R}_{\mathfrak{g}}$ into a *-subalgebra of $\mathcal{A}_{\Omega}$ could simplify this problem, as suggested by Exa.
+9.3, is not yet clarified.
+We finish this paper with a map whose continuity is more subtle than that of the inference.
 Example 9.5 (Entropy distance I).
-Let $\tau:=\mathds{1}_{n}/n$ and let $\mathcal{E}(\mathcal{R}):=\mathcal{E}_{\tau}(\mathcal{R})$ be an exponential family, see Ex. 9.3. The entropy distance from $\mathcal{E}(\mathcal{R})$ is
-$d:\mathcal{D}(\mathrm{M}_{n})\to\mathbb{R}\,,\quad\rho_{1}\mapsto\inf_{\rho_{2}\in\mathcal{E}(\mathcal{R})}S(\rho_{1},\rho_{2})\,.$ (9.6)
-Using the von Neumann entropy $S=\phi_{\tau}+\log(n)$, we define
-$\Psi:\mathcal{D}(\mathcal{R})\to\mathcal{D}(\mathrm{M}_{n})\,,\quad\sigma\mapsto\operatorname{argmax}_{\rho\in\pi^{-1}(\sigma)}S(\rho)\,.$
-Clearly, the map $\Psi$ equals the inference map $\Psi_{\tau}$ defined earlier in Ex. 9.1. The entropy distance can be written as the difference
+The entropy distance from the exponential family $\mathcal{E}(\mathcal{R})$ is defined by
+$d:\mathcal{D}(\mathrm{M}_{n})\to\mathbb{R}\,,\quad\rho_{1}\mapsto\inf_{\rho_{2}\in\mathcal{E}(\mathcal{R})}S(\rho_{1},\rho_{2})$ (9.3)
+and equals the difference
 $d(\rho)=S(\Psi\circ\pi(\rho))-S(\rho)\,,\quad\rho\in\mathcal{D}(\mathrm{M}_{n})$ (9.4)
-between the value of the von Neumann entropy at $\rho$ and its maximal value on the fiber of $\pi:\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathcal{R})$ that contains $\rho$, see p. 1288 in [52].
-Formula (9.7) suggests studying the rI-projection [19]
-$\Pi:\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathrm{M}_{n})\,,\quad\Pi=\Psi\circ\pi\,.$
-By (9.3), the range of $\Pi$ is the rI-closure of $\mathcal{E}(\mathcal{R})$. We also consider the restriction $d|_{\mathcal{D}(\mathcal{A})}$ of $d$ and the following restriction of $\Pi$.
-$\Pi|_{\mathcal{D}(\mathcal{A})}:\mathcal{D}(\mathcal{A})\to\mathcal{D}(\mathcal{A})\,,\quad\rho\mapsto\Pi(\rho)$
-Here, $\mathcal{A}$ is real *-subalgebra of $\mathrm{M}_{n}$ that includes $\mathcal{R}$.
-The map $\Pi|_{\mathcal{D}(\mathcal{A})}$ is well defined because $\Psi(\mathcal{D}(\mathcal{R}))$ is included in $\mathcal{D}(\mathcal{A})$, as observed in Ex.
-9.3 a).
-Example 9.6 (Entropy distance II).
-The exponential family $\mathcal{E}(\mathcal{R}_{\mathfrak{g}})$ of local Hamiltonians (Ex.
-9.4) is interesting because the entropy distance from it quantifies many-body correlations.
-Amari [4] and Ay [6] studied this type of correlation measure in probability theory.
-Linden et al.
-[33] introduced it as a difference of von Neumann entropies like formula (9.7), see also [58, Sec.
-1.4.2].
-And Zhou [59] proved the equivalence of the two representations (9.6) and (9.7) for density matrices of maximal rank $n$, see also [38]; the equivalence is true without rank restrictions as presented in Ex.
-9.5.
-We quote from [52, Lemma 5.15 and 4.5] under the assumptions of Ex. 9.5.
+between the value of the von Neumann entropy at $\rho$ and the maximal value on the fiber of $\pi$ that contains $\rho$, see p.
+1288 in [53].
+Formula (9.4) suggests studying the continuity of $d$ through the rI-projection
+$\Pi:\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathrm{M}_{n})\,,\quad\Pi:=\Psi\circ\pi\,.$
+As per Def.
+5.2 and equation (D8) in [53], the density matrix $\Pi(\rho_{1})$ is the generalized rI-projection of $\rho_{1}\in\mathcal{D}(\mathrm{M}_{n})$ to $\mathcal{E}(\mathcal{R})$, which is a well-known concept in probability theory [18], and which is defined as follows.
+A sequence $(\tau_{i})\subset\mathcal{D}(\mathrm{M}_{n})$ rI-converges to $\rho_{2}\in\mathcal{D}(\mathrm{M}_{n})$ if $\lim_{i}S(\rho_{2},\tau_{i})=0$ holds.
+If every sequence $(\tau_{i})\subset\mathcal{E}(\mathcal{R})$ satisfying $\lim_{i}S(\rho_{1},\tau_{i})=d(\rho_{1})$ rI-converges, independently of the sequence, to a unique $\rho_{2}\in\mathcal{D}(\mathrm{M}_{n})$, not necessarily in $\mathcal{E}(\mathcal{R})$, then $\rho_{2}$ is the generalized rI-projection of $\rho_{1}$ to $\mathcal{E}(\mathcal{R})$.
+We quote from Lemma 5.15 and Lemma 4.5 in [53].
 Lemma 9.6.
-Le $\mathcal{A}$ be real *-subalgebra of $\mathrm{M}_{n}$ that includes $\mathcal{R}$.
   1. a)
-    For every $\rho\in\mathcal{D}(\mathcal{A})$, the map $\Pi|_{\mathcal{D}(\mathcal{A})}$ is continuous at $\rho$ if and only if $d|_{\mathcal{D}(\mathcal{A})}$ is continuous at $\rho$.
+    For every $\rho\in\mathcal{D}(\mathrm{M}_{n})$ the rI-projection $\Pi$ is continuous at $\rho$ if and only if the entropy distance $d$ is continuous at $\rho$.
   2. b)
     For every $\sigma\in\mathcal{D}(\mathcal{R})$, the inference map $\Psi$ is continuous at $\sigma$ if and only if $d$ is continuous at every point in the fiber $\pi^{-1}(\sigma)$.
-With reference to Ex. 9.5, we simplify the continuity problem of the entropy distance (and of other invariant maps) using *-subalgebras.
-Remark 9.8 (Continuity of the rI-projection).
-Let $\mathcal{A}$ be a *-subalgebra of $\mathrm{M}_{n}$ such that $\mathcal{R}\subset\mathcal{A}$.
-As in Rem.
-8.2, the map $\pi$ factors into $\pi=\pi_{\mathcal{R}}\circ\pi_{\mathcal{A}}$.
-Then, for every $\rho\in\mathcal{D}(\mathrm{M}_{n})$ the rI-projection $\Pi$ is continuous at $\rho$ if and only if $\Pi|_{\mathcal{D}(\mathcal{A})}$ is continuous at $\pi_{\mathcal{A}}(\rho)$.
-This assertions needs a justification.
-By its definition, the map $\Pi=\Psi\circ\pi$ is invariant under $\pi_{\mathcal{A}}$ in the sense that $\Pi=\Pi\circ\pi_{\mathcal{A}}$ holds.
-This shows $\Pi=\Pi|_{\mathcal{D}(\mathcal{A})}\circ\pi_{\mathcal{A}}$.
-The claim then follows from Lemma 8.1 b), by letting $\mathcal{R}_{1}:=\mathrm{M}_{n}$ and $\mathcal{R}_{2}:=\mathcal{A}$, and by taking $\Pi|_{\mathcal{D}(\mathcal{A})}$ as the map $f:\mathcal{D}(\mathcal{R}_{2})\to T$.
-The assumptions of the lemma are met since $\pi_{\mathcal{A}}$ is open by Thm.
-1.2.
+Example 9.7 (Entropy distance II).
+The entropy distance from the exponential family $\mathcal{E}(\mathcal{R}_{\mathfrak{g}})$ of local Hamiltonians (Ex.
+9.4) is interesting because it quantifies many-body correlations.
+Amari [4] and Ay [6] studied this type of correlation measures in probability theory.
+Linden et al.
+[33] introduced it to quantum mechanics as a difference of von Neumann entropies like formula (9.4), see also [59, Sec.
+1.4.2].
+Zhou [60] proved the equality of the two representations (9.3) and (9.4) for density matrices of maximal rank $n$, see also [38]; the equality is true without rank restrictions as stated in Ex.
+9.5.
+Proposition 9.8.
+For every $\rho\in\mathcal{D}(\mathrm{M}_{n})$, the entropy distance $d$ is continuous at $\rho$ if and only if its restriction $d|_{\mathcal{D}(\mathcal{A})}$ is continuous at $\pi_{\mathcal{A}}(\rho)$.
+Proof.
+Let $\rho\in\mathcal{D}(\mathrm{M}_{n})$. By Lemma 9.6 a), the entropy distance $d$ is continuous at $\rho$ if and only if the rI-projection $\Pi$ is continuous at $\rho$. The inference map
+$\Psi^{\mathcal{A}}:\mathcal{D}(\mathcal{R})\to\mathcal{D}(\mathcal{A})\,,\quad\sigma\mapsto\operatorname{argmax}_{\eta\in\pi_{\mathcal{R}}^{-1}(\sigma)}S(\eta)$
+has the same values as $\Psi$, whose image $\Psi(\mathcal{D}(\mathcal{R}))$ is included in $\mathcal{D}(\mathcal{A})$ by Ex. 9.3 a). Therefore,
+$\Pi=\Psi\circ\pi=\Psi\circ\pi_{\mathcal{R}}\circ\pi_{\mathcal{A}}$
+is continuous at $\rho$ if and only if $\Psi^{\mathcal{A}}\circ\pi_{\mathcal{R}}\circ\pi_{\mathcal{A}}$ is continuous at $\rho$, if and only if $\Pi^{\mathcal{A}}\circ\pi_{\mathcal{A}}$ is continuous at $\rho$, where
+$\Pi^{\mathcal{A}}:\mathcal{D}(\mathcal{A})\to\mathcal{D}(\mathcal{A})\,,\quad\Pi^{\mathcal{A}}:=\Psi^{\mathcal{A}}\circ\pi_{\mathcal{R}}\,.$
+Since $\pi_{\mathcal{A}}$ is open by Thm.
+1.2, it follows from Lemma 8.1 b) that $\Pi^{\mathcal{A}}\circ\pi_{\mathcal{A}}$ is continuous at $\rho$ if and only if $\Pi^{\mathcal{A}}$ is continuous at $\pi_{\mathcal{A}}(\rho)$.
+By equation (9.4),
+$d|_{\mathcal{D}(\mathcal{A})}(\eta)=S(\Psi^{\mathcal{A}}\circ\pi_{\mathcal{R}}(\eta))-S(\eta)\,,\quad\eta\in\mathcal{D}(\mathcal{A})$
+holds.
+Hence, $\Pi^{\mathcal{A}}$ is continuous at $\pi_{\mathcal{A}}(\rho)$ if and only if $d|_{\mathcal{D}(\mathcal{A})}$ is continuous at $\pi_{\mathcal{A}}(\rho)$, again by Lemma 5.15 1) in [53].
+∎
 Example 9.9 (Entropy distance III).
-Rem.
-9.8 helps solve the continuity problem of the entropy distance from $\mathcal{E}(\mathcal{R})$ for $\mathcal{R}:=\operatorname{span}_{\mathbb{R}}(\mathds{1}_{3},X\oplus 1,Z\oplus 0)$, based on a solution for $d|_{\mathcal{A}(\mathcal{D})}$ in the algebra $\mathcal{A}:=\mathrm{M}_{2}(\mathbb{R})\oplus\mathrm{M}_{1}(\mathbb{R})$ that was obtained using asymptotic curvature estimates [52].
-Real *-subalgebras are excluded from Rem. 9.8 but the conclusion is still true, as Ex. 7.2 can replace Thm. 1.2 in this remark.
+Prop.
+9.8 helps solve the continuity problem of the entropy distance from $\mathcal{E}(\mathcal{R})$ for $\mathcal{R}:=\operatorname{span}_{\mathbb{R}}(\mathds{1}_{3},X\oplus 1,Z\oplus 0)$ using the solution of the continuity problem of the restriction $d|_{\mathcal{D}(\mathcal{A})}$ to the real *-subalgebra $\mathcal{A}:=\mathrm{M}_{2}(\mathbb{R})\oplus\mathrm{M}_{1}(\mathbb{R})$ of $\mathrm{M}_{3}$.
+This solution was obtained from asymptotic curvature estimates [53].
+Real *-subalgebras are excluded from Proposition 9.8 but the conclusion is still true, as Ex.
+7.2 can replace Thm.
+1.2 in the proof of Prop.
+9.8.
 We recall from Ex.
@@
 $\mathcal{G}_{d}:=\big{[}\ket{+}\!\!\bra{+}\oplus 0,M\big{)}=\big{\{}(1-\lambda)\ket{+}\!\!\bra{+}\oplus\lambda\colon\lambda\in[0,\frac{1}{2})\big{\}}$
-and continuous at every point in the complement.
-(Unitary similarity with respect to $U\oplus 1$ matches $\mathcal{R}$ with the problem considered in [52], as the similarity with respect to $U:=\exp(\mathrm{i}\frac{\pi}{3\sqrt{3}}(X+Y+Z))$ permutes the Pauli matrices $X,Y,Z$ cyclicly.) Lemma 9.7 then shows that the restricted rI-projection $\Pi|_{\mathcal{D}(\mathcal{A})}$ is discontinuous at every point in $\mathcal{G}_{d}$ and continuous at every point in the complement.
-Now Rem.
-9.8 proves that the rI-projection $\Pi$ is discontinuous at every point in $\pi_{\mathcal{A}}^{-1}(\mathcal{G}_{d})$ and continuous at every point in the complement.
-A second application of Lemma 9.7 finally shows that the entropy distance $d$ is discontinuous at every point in $\pi_{\mathcal{A}}^{-1}(\mathcal{G}_{d})$ and continuous at every point in the complement.
+and continuous at every point in the complement [^6] Unitary similarity with respect to $U:=\exp(\mathrm{i}\frac{\pi}{3\sqrt{3}}(X+Y+Z))$ permutes the Pauli matrices cyclicly.
+With respect to $U\oplus 1$ it matches $\mathcal{R}$ and the problem considered in [53]..
+Prop.
+9.8 then shows that the entropy distance $d$ is discontinuous at every point in $\pi_{\mathcal{A}}^{-1}(\mathcal{G}_{d})$ and continuous at every point in the complement.
+This result is consistent with the assertion of Ex.
+9.3 c) that the inference map $\Psi$ is discontinuous at $M$ and continuous anywhere else in the ellipse $\mathcal{D}(\mathcal{R})$, as required by Lemma 9.6 b).
 The points in the set $\pi_{\mathcal{A}}^{-1}(\mathcal{G}_{d})$ are explicitly described in Ex.
 8.3 b).
-The result is consistent with the inference $\Psi$ being discontinuous at $M$ and continuous at every point different from $M$ (see Ex.
-9.3 c)) as per Lemma 9.7 b).
 Acknowledgments.
-The author is grateful to several people for discussions related to this work, among them are Achim Clausing, Chi-Kwong Li, Tim Netzer, Ivan Todorov, and Ilya M.
-Spitkovsky.
+The author is grateful to Ariel Caticha, Achim Clausing, Chi-Kwong Li, Tim Netzer, Ilya M.
+Spitkovsky, and Ivan Todorov for remarks, feedback, and discussions.
 The author gave a lecture on the results of this paper in the special session “Numerical Ranges” of the conference IWOTA 2024 in Canterbury, UK, and wishes to thank the organizers for this opportunity.